### v12

# Defining implication relation for classical logic

Abstract
In classical logic, “P implies Q” is equivalent to “not-P or Q”. It is well known that the equivalence is problematic. Actually, from “P implies Q”, “not-P or Q” can be inferred (“Implication-to-Disjunction” is valid), whereas from “not-P or Q”, “P implies Q” cannot be inferred in general (“Disjunction-to-Implication” is not generally valid), so the equivalence between them is invalid in general. This work aims to remove the incorrect Disjunction-to-Implication from classical logic (CL). The logical system (the logic IRL) this paper proposes has the expected properties: (a) CL is obtained by adding Disjunction-to-Implication to IRL, and (b) Disjunction-to-Implication is not derivable in IRL; while (c) fundamental laws in classical logic, including law of excluded middle (LEM) and principle of double negation, law of non-contradiction (LNC) and ex contradictione quodlibet (ECQ), conjunction elimination and disjunction introduction, and hypothetical syllogism and disjunctive syllogism, are all retained in IRL.
Keywords implication, conditional, material implication, paradox, classical logic, propositional logic, first-order logic, proof system, Hilbert system, natural deduction, semantics, algebraic logic
1 Introduction
In classical logic, “$P$ implies $Q$” ($P\to Q$) is taken to be logically equivalent to “it cannot be $P$ and not $Q$” ($\neg(P\wedge\neg Q)$) or “not-$P$ or $Q$” ($\neg P\vee Q$) by duality. This equivalence can be viewed as the definition of implication by disjunction and negation ($P\to Q=\neg P\vee Q$, called material implication), or used to define disjunction by implication and negation, so $P\to Q$ and $\neg P\vee Q$ can be replaced with one another in a derivation.
It is well known that this definition of implication is problematic in that it leads to counter-intuitive results or “paradoxes”. There are many paradoxes of material implication (see e.g. Bronstein, 1936). For instance, the theorem
$(P\to Q)\vee(Q\to P)$
of classical logic means that any two propositions must be related by implication (total order). This does not agree with a large number of situations in the real world. For example, let $P=$ “Alice is in Athens” and $Q=$ “Alice is in London” (given an instant in the future), then both $P\to Q$ and $Q\to P$ cannot be true, so $(P\to Q)\vee(Q\to P)$ cannot be true.
Many efforts have been made to resolve this problem and important non-classical logics are developed, such as modal logics which introduce the “strict implication”, relevance logics which require that the antecedent and consequent of implications to be relevantly related (so they are paraconsistent logics which reject ex contradictione quodlibet (ECQ)).
However, classical logic exists on its own reason (see e.g. Fulda, 1989). And implication is the kernel concept in logic: “implies” means the same as that “of ordinary inference and proof” (Lewis, 1912). So it is not satisfactory that the uniquely fundamental, simple and useful classical logic has a wrong definition of implication – the material implication.
The paper explains why “Implication-to-Disjunction” ($(P\to Q)\to\neg P\vee Q$) is valid whereas “Disjunction-to-Implication” ($\neg P\vee Q\to(P\to Q)$) is not, by describing firstly the original meaning of “implication”. So the goal of this work is a logic in which Disjunction-to-Implication is removed while fundamental laws in classical logic such as law of excluded middle (LEM) and principle of double negation, law of non-contradiction (LNC) and ex contradictione quodlibet (ECQ), conjunction elimination and disjunction introduction, and hypothetical syllogism (transitivity of implication) and disjunctive syllogism, are all retained.
2 Original meaning of implication
2.1 Four situations
Let $P$ and $Q$ be propositions. Consider how $P$ affects $Q$. There are four possible situations about $Q$ when $P$ is true:
  1. 1.
    when $P$ is true, $Q$ is necessarily true;
  2. 2.
    when $P$ is true, $Q$ is coincidentally true;
  3. 3.
    when $P$ is true, $Q$ is necessarily false;
  4. 4.
    when $P$ is true, $Q$ is coincidentally false.
Where:
  • “necessarily” means “with a certain mechanism”;
  • “coincidentally” means “without a certain mechanism”.
And “certain mechanism” means one of the two:
  1. 1.
    It is certain literally or physically (in some scope, cf. Subsection 4.6), for example, “in the open air” is a physical mechanism of “if it rains, then the ground gets wet” (cf. Example 4.6.2);
  2. 2.
    It is certain formally or logically, for example, “from $(P\vee Q)\wedge\neg P$ to $Q$” is a logical mechanism.
2.2 Original meaning of implication
The original meaning of implication $P\to Q$ is described as follows.
  • $P\to Q$ is true, if and only if whenever $P$ is true, $Q$ is necessarily true, no matter whether $Q$ is true or false when $P$ is false. Equivalently,
  • $P\to Q$ is false, if and only if whenever $P$ is true, $Q$ is not necessarily true (i.e. $Q$ is coincidentally true, necessarily false, or coincidentally false), no matter whether $Q$ is true or false when $P$ is false.
According to this description, implication from propositions $P$ to $Q$ is (can be defined as) a relation determined by a certain mechanism (physical or logical) from $P$ to $Q$, which guarantees that $Q$ is true whenever $P$ is true. Without such a mechanism, even if $Q$ is true when $P$ is true, $P\to Q$ is still false since $Q$ is just coincidentally true when $P$ is true.
The original meaning of bi-implication $P\leftrightarrow Q$ is defined similarly as that “$P\leftrightarrow Q$ is true” means “there is a certain mechanism between $P$ and $Q$ in some scope, which guarantees bi-directional truth-preserving”.
2.3 Non-truth-functionality of implication
[…]
Let $L$ be the set of all the formulas representing a propositional language. Let $U$ be the set of all valuation functions from $L$ to the set of truth values $\{\mathrm{T,F}\}$. Then:
According to the meaning of conjunction, there exists a function $f:\{\mathrm{T,F}\}^{2}\to\{\mathrm{T,F}\}$ such that for any $v\in U$ and any $\phi,\psi\in L$, $v(\phi\wedge\psi)=f(v(\phi),v(\psi))$, That is, $\wedge$ is truth-functional.
In contrast, according to the original meaning of implication, there exists no function $f:\{\mathrm{T,F}\}^{2}\to\{\mathrm{T,F}\}$ such that for any $v\in U$ and any $\phi,\psi\in L$, $v(\phi\to\psi)=f(v(\phi),v(\psi))$. That is, $\to$ is not truth-functional.
For instance, in terms of the meaning of conjunction, if we know that $P$ is true and $Q$ is true, then we know for sure that $P\wedge Q$ is true; whereas, in terms of the original meaning of implication, even if we know that $P$ is true and $Q$ is true, we are not sure about whether $P\to Q$ is true or false, because this depends on whether there is “a certain mechanism” from $P$ to $Q$ to guarantee that whenever $P$ is true $Q$ must be true.
In fact, we know surely the truth value of $P\wedge Q$ for all possible four combinations of truth values $\{\mathrm{TT,TF,FT,FF}\}$ of $P$ and $Q$, whereas we do not know the truth value of $P\to Q$ from that of $P$ and $Q$ except for the only one combination $\mathrm{TF}$: only when $P$ is true and $Q$ is false, we know for sure that $P\to Q$ is false, because (only) $P\wedge\neg Q\to\neg(P\to Q)$ is valid according to the original meaning of implication (cf. Subsection 2.4).
From the semantic viewpoint, implication $\to$ is a binary relation rather than a binary operation like conjunction $\wedge$ and disjunction $\vee$. Specifically, the implication $\to$, as a binary relation on $L$, the truth value of $\phi\to\psi$ is determined by whether $(\phi,\psi)$ belongs to the set of the relation $\to$ that is a subset of $L\times L$ (semantically, it is just the T-set of the implication, cf. Subsection 4.1), rather than by the truth value of $\phi$ and the truth value of $\psi$. We call $\to$ implication relation when emphasizing this fact.
So is for bi-implication $\leftrightarrow$.
2.4 Incorrectness of “material implication” as implication
Consider the relation between $P\to Q$ and $\neg P\vee Q$ in terms of the semantics of implication defined above.
$\neg P\vee Q$ is true” consists of three cases: $P$ is true and $Q$ is true, $P$ is false and $Q$ is true, $P$ is false and $Q$ is false. The first case “$P$ is true and $Q$ is true” includes not only the sole situation where $P\to Q$ is true (i.e. $Q$ is necessarily true when $P$ is true), but also the the situation “$P$ is true and $Q$ is coincidentally true” which is not a context where $P\to Q$ is true. In other words, “$\neg P\vee Q$ is true” contains not only the case where “$P\to Q$ is true” but also more cases. In symbols, $(P\to Q)\to\neg P\vee Q$ holds whereas $\neg P\vee Q\to(P\to Q)$ does not.
$\neg P\vee Q$ is false” contains only one case: $P$ is true and $Q$ is false that includes two sub-cases, i.e. “$P$ is true and $Q$ is necessarily false” and “$P$ is true and $Q$ is coincidentally false”. The two sub-cases is a subset of situations where $P\to Q$ is false (cf. Subsection 2.2). In symbols, ¬($\neg(\neg P\vee Q)\to\neg(P\to Q)$ (i.e. $P\wedge\neg Q\to\neg(P\to Q)$) holds, whereas $\neg(P\to Q)\to\neg(\neg P\vee Q)$ does not. This result is the same as above in terms of contraposition.
In summary, $(P\to Q)\leftrightarrow\neg P\vee Q$ is generally not a tautology since
$(P\to Q)\to\neg P\vee Q$ (Implication-to-Disjunction) is valid while
$\neg P\vee Q\to(P\to Q)$ (Disjunction-to-Implication) is generally invalid.
Some researchers have addressed this issue with different motivations or explanations, such as MacColl (1880), Bronstein (1936), Woods (1967), and Dale (1974).
2.5 Difference from “strict implication”
The meaning of “strict implication” of modal logics is NOT the same as the original meaning of implication described above. In a classical modal logic:
  • $\Box(P\to Q)$ is true means “necessarily $P$ implies $Q$” is true. Equivalently,
  • $\Box(P\to Q)$ is false, or $\neg\Box(P\to Q)$ is true, or $\Diamond\neg(P\to Q)$ is true, means “not necessarily $P$ implies $Q$” is true, or “possibly not $P$ implies $Q$” is true.
The negation of strict implication, “it is possible not that $P$ implies $Q$”, is not excluding the situation “it is possible that $P$ implies $Q$”. This is different from the original meaning of implication defined in previous Subsection 2.2.
As a matter of fact, “strict implication” does not define the meaning of implication itself: let $S=P\to Q$, the strict implication $\Box(P\to Q)=\Box S$ says just “necessarily $S$”, nothing about the implication $P\to Q$ itself.
2.6 Construction of the right logic
The goals of this work in construction of the logic are:
  1. 1.
    Classical scope. The logical system to be constructed is “in the classical scope”. For example, it is a two-valued logic and a sub-system of classical logic, so that any operators new to classical logic like modal operators are not considered.
  2. 2.
    Not conflicting with the original meaning of implication. The system should not contain invalid theorems in terms of the meaning of operators, especially the original meaning of implication $\to$. For example, Disjunction-to-Implication $\neg\phi\vee\psi\to(\phi\to\psi)$ must be rejected.
  3. 3.
    Strength. Under the above two conditions, the system should be as strong as possible. For example, both LEM and ECQ are not rejected. LEM is irrelevant to the original meaning of implication. ECQ, which is necessary for retaining both disjunction introduction and disjunctive syllogism (cf. Section 9), does not conflict with the original meaning of implication (cf. Subsection 4.1). And Implication-to-Disjunction $(\phi\to\psi)\to\neg\phi\vee\psi$ is kept since it follows directly from the original meaning of implication as explained previously.
There are many choices of ways for constructing logics. Two main approaches are:
  1. 1.
    Choose bi-implication $\leftrightarrow$ as the primitive one, then define implication $\to$ by bi-implication with the rule $\phi\to\psi\Leftrightarrow\phi\wedge\psi\leftrightarrow\phi$ or $\phi\to\psi\Leftrightarrow\phi\vee\psi\leftrightarrow\psi$.
  2. 2.
    Choose implication $\to$ as the primitive one, then define bi-implication $\leftrightarrow$ by implication with the rule $\phi\leftrightarrow\psi\Leftrightarrow(\phi\to\psi)\wedge(\psi\to\phi)$.
Because of its symmetry, bi-implication-based approach has an advantage in structure. The logic IRL presented in Section 3 and its algebraic system IRLA in Section 4, are constructed in the bi-implication-based approach. The natural deduction for IRL is constructed in the implication-based approach that is more suitable to it.
3 Propositional logic IRL
IRL (Implication-Relation Logic) is used to refer the logic presented in this section.
3.1 Definitions and notations
The following definitions and notations are adopted for the systems proposed in this paper. The description of concepts is general, e.g. “formula” in this subsection may mean not only the “(well-formed) formula” of a propositional language or first-order language, but also a statement like a rule expression or a sentence of some meta-language being used.
Definitions
Definition 3.1.1 (Rule of inference).
Let $X$ be a (normally non-empty) finite set of formulas (premises) and $y$ be a formula (conclusion).
A (single-conclusion) rule of inference is an expression of the form
$X\Rightarrow y$
that is interpreted as that if every element of $X$ holds, then $y$ holds.
Let $Y=\{y_{1},...,y_{n}\}$ be a non-empty finite set of formulas (conclusions).
A multiple-conclusion rule
$X\Rightarrow Y$
is defined as the set of single-conclusion rules $\{X\Rightarrow y_{1},...,X\Rightarrow y_{n}\}$.
A bi-inference rule
$X\Leftrightarrow Y$
is defined as the union of $X\Rightarrow Y$ and $Y\Rightarrow X$ (treat a single-conclusion rule as a singleton for the union operation).
Remark 3.1.1.
Empty-premise rule: In Definition 3.1.1, a rule with an empty set of premises represents actually axiom(s) or theorem(s).
Domain-specific rule: A domain-specific rule is sometimes called an axiom or a theorem in that domain. For example, in a Boolean lattice extended from a Boolean algebra, $x\leq y\Leftrightarrow x\wedge y=x$ is usually listed as the defining axiom for its partial order ($\leq$), although it is essentially a (bi-inference) domain-specific rule.
Definition 3.1.2 (Proof and theorem).
Let $X$ be a (possibly empty) finite set of formulas (premises) and $y$ be a formula (conclusion).
A proof of $y$ from $X$, denoted
$X\vdash y,$
is a non-empty finite sequence of formulas $(y_{1},...,y_{n})$ with $y_{n}=y$, where $y_{k}\;(1\leq k\leq n)$ is an axiom, a theorem, or a result from a subset of $X\cup\{y_{1},...,y_{k-1}\}$ by applying a rule of inference.
  • An unconditional proof is a proof with an empty set of premises. Formula $y$ is called a theorem if there exists an unconditional proof of $y$, denoted
    $\vdash y.$
  • A Conditional proof is a proof with a non-empty set of premises, i.e. $X\vdash y$ where $X$ is non-empty. Note that a member of the set $X$ of premises is not a member of its proof sequence $(y_{1},...,y_{n})$.
The length or steps of a proof is the length of its proof sequence.
Remark 3.1.2.
According to Definition 3.1.2:
  • All axioms are theorems. Therefore, any member of a proof sequence is just either a theorem or a result from a subset of premises and previous members of the proof sequence by applying a rule.
  • In a proof sequence $(y_{1},...,y_{n})$ of an unconditional proof, each subsequence of the form $(y_{1},...,y_{k})\,(1\leq k\leq n)$ is a unconditional proof, thus every member $y_{k}\,(1\leq k\leq n)$ is a theorem.
Definition 3.1.3 (Derivable rule).
Let $X$ and $Y$ be finite sets of formulas where $Y$ is non-empty.
A rule $X\Rightarrow Y$ is derivable if and only if there exists a proof $X\vdash Y$ where rules used are only existing ones. An existing rule is a primitive rule or a derived rule.
In terms of the definitions of proof and derivable rule, any primitive rule is derivable.
Remark 3.1.3 (Meta rule and meta proof).
If a premise or a conclusion of a rule or a proof contains a symbol not belonging to the object language, then the rule or the proof is a meta rule or meta proof. For example, the rule $(\phi\vdash\psi)\Rightarrow(\vdash\phi\to\psi)$ of conditional proof is a meta rule, a proof $(\phi\vdash\psi)\vdash(\vdash\phi\to\psi)$ of the deduction theorem is a meta proof.
Definition 3.1.4 (Uniform substitution).
Let $x_{1},...,x_{k}$ be variables, $t_{1},...,t_{k}$ be terms, and $E$ be an expression containing variables.
A uniform substitution is a set $\{t_{1}/x_{1},...,t_{k}/x_{k}\}$ and $E\{t_{1}/x_{1},...,t_{k}/x_{k}\}$ means to (simultaneously) substitute $t_{i}$ for every occurrence of each $x_{i}$ in $E$. A member $x_{i}/x_{i}$ in a substitution $\{...,x_{i}/x_{i},...\}$ can be omitted, e.g. $\{t_{1}/x_{1},x_{2}/x_{2},t_{3}/x_{3}\}$ is the same as $\{t_{1}/x_{1},t_{3}/x_{3}\}$.
The rule of uniform substitution is the form $E\Rightarrow E\{t_{1}/x_{1},...,t_{k}/x_{k}\}$, where $E$ is an axiom, a theorem, a rule, or a meta rule of a formal system.
The rule of uniform substitution is common to logical systems for instantiating (schemas of) axioms, theorems, rules and meta rules, and it is often used tacitly.
Definition 3.1.5 (Subformula replacement).
Let $x,y,z$ be formulas.
The expression $z[x\mapsto y]$ denotes any member of the set of all possible results where each result is obtained via replacing one or more arbitrarily selected occurrence(s) of $x$ by $y$ in $z$.
For instance, $(x\wedge y\to x\wedge y)[x\wedge y\mapsto y]$ represents any result in the set $\{y\to x\wedge y,\,x\wedge y\to y,\,y\to y\}$.
Presentation of proofs
The presentation format of a proof is as follows.
An unconditional proof $\vdash\psi$ of length $n$ is presented in the format:
$1$ $\psi_{1}$ reason
$\vdots$ $\vdots$ $\vdots$
$n-1$ $\psi_{n-1}$ reason
$n$ $\psi$ reason
A conditional proof $\{\phi_{1},...,\phi_{m}\}\vdash\psi$ of length $n$ is presented in the format:
$1$ $\phi_{1}$ Premise (Assumption)
$\vdots$ $\vdots$ $\vdots$
$m$ $\phi_{m}$ Premise (Assumption)
$m+1$ $\psi_{1}$ reason
$\vdots$ $\vdots$ $\vdots$
$m+n-1$ $\psi_{n-1}$ reason
$m+n$ $\psi$ reason
A proof (unconditional or conditional) can has other proofs (unconditional or conditional) embedded within it. In the presentations, this nested structure of proofs must be clearly indicated in a consistent manner such as indentation used in this paper.
An example that an unconditional proof has an inner conditional proof:
1 $\psi_{11}$ theorem-name
2     $\phi_{21}$ Premise
3     $\psi_{21}$ theorem-name
4     $\psi_{22}$ 2,3: rule-name (From $\{2,3\}$ by rule-name)
5 $\phi_{21}\to\psi_{22}$ 2–4: Rule of conditional proof
6 $\psi_{12}$ 1,5: rule-name
Note that the result of a conditional proof (by the rule of conditional proof) is not a member of its own proof sequence, so it must be unindented to just one level higher (one layer outer) as shown in the example (line 5).
[…]
Hence, in the presentation of a proof, the format
$\alpha_{0}$
$1$     $\backsim\alpha_{1}$
$\vdots$     $\vdots$
$n$     $\backsim\alpha_{n}$
is a shorthand of the format
$1$ $\alpha_{0}\backsim\alpha_{1}$
$\vdots$ $\vdots$
$n$ $\alpha_{n-1}\backsim\alpha_{n}$
A note on nested proofs
By definition (Definition 3.1.2, and cf. Remark 3.1.2), any member of a proof sequence must be either a theorem or a result from a subset of premises and previous members by a rule. Specifically, let $X\vdash y_{n}$ has a proof sequence $(y_{1},...,y_{n})$, then any $y_{k}$ must be in one of the following two cases:
  • $y_{k}$ is a theorem, or it is assumed be a theorem in a premise of an outer proof, or it is a theorem derived in an outer proof;
  • $y_{k}$ is obtained from a subset of $X\cup\{y_{1},...,y_{k-1}\}$ by applying a rule, or by an assumed rule in a premise of an outer proof, or by a rule derived in an outer proof.
No other cases are allowed.
This is further illustrated in Table 1.
Attention must be paid to this, especially for a nested conditional proof.
Table 1: Only two cases allowed for members of proof sequence
An outer proof The inner proof Comment
$A_{1}$
$\vdots$
$A_{m}$$\{A_{1},...,A_{m}\}$ is the set of premises/assumptions of the outer proof a
$B_{1}$
$\vdots$
$B_{i}$ For $B_{i}$, a member of the outer proof sequence, only two cases are allowed: it is a theorem, or it is a result from a subset of $\{A_{1},...,A_{m}\}\cup\{B_{1},...,B_{i-1}\}$ by applying a rule b
$\vdots$
$B_{j}$
$P_{1}$
$\vdots$
$P_{n}$$\{P_{1},...,P_{n}\}$ is the set of premises/assumptions of the inner proof a
$Q_{1}$
$\vdots$
$Q_{k}$ For $Q_{k}$, a member of the inner proof sequence, only two cases are allowed: (a) it is a theorem, or it is assumed be a theorem in $\{A_{1},...,A_{m}\}$, or it is a theorem in $\{B_{1},...,B_{j}\}$; or (b) it is a result from a subset of $\{P_{1},...,P_{n}\}\cup\{Q_{1},...,Q_{k-1}\}$ by a rule, or by a rule assumed in $\{A_{1},...,A_{m}\}$, or by a rule in $\{B_{1},...,B_{j}\}$
$\vdots$
  • a
    If the set of premises is empty, it is an unconditional proof; otherwise, it is a conditional proof.
  • b
    If the outer proof itself is embedded in another proof, treat it the same way as the inner proof in the table.
Example 3.1.1 (An illegal proof).
Proof of $\phi\to(\psi\to\phi)$:
1     $\phi$ Premise (outer) 2       $\psi$ Premise (inner) 3       $\phi$ 1: Reiteration – Not allowed 4     $\psi\to\phi$ 2–3: Rule of conditional proof 5 $\phi\to(\psi\to\phi)$ 1–4: Rule of conditional proof
Line 3 is not allowed because $\phi$ in this line does not belong to any of the allowed cases: (a) it is a theorem or a theorem indicated in an outer proof ($\phi$ is not assumed be a theorem in the outer premise i.e. line 1); (b) it is a result from a subset of premises and previous members of its own layer (that must be a subset of $\{\psi\}$ from line 2) by applying a rule or by an rule in an outer proof.
Example 3.1.2 (A legal proof).
Proof of $(\top\to\phi)\to(\psi\to\phi)$:
1     $\top\to\phi$ Premise (outer) 2       $\psi$ Premise (inner) 3       $\top$ A theorem ($\top$ is indeed a theorem) 4     $\psi\to\top$ 2–3: Rule of conditional proof 5     $\psi\to\phi$ 4,1: Rule of transitivity 6 $(\top\to\phi)\to(\psi\to\phi)$ 1–5: Rule of conditional proof
Sometimes, a meta proof (for a meta language) contains an ordinary proof (for an object language) such as the following example, which details a meta proof of a meta rule.
Example 3.1.3 (A legal meta proof).
Proof of $(\vdash\psi)\Rightarrow(\vdash\phi\to\psi)$:
1     $\vdash\psi$ Premise (meta proof) 2       $\phi$ Premise (proof) 3       $\psi$ 1 ($\psi$ is assumed in line 1 be a theorem) 4     $\phi\vdash\psi$ 2–3: Definition of conditional proof 5     $\vdash\phi\to\psi$ 4: Rule of conditional proof 6 $(\vdash\psi)\vdash(\vdash\phi\to\psi)$ 1–5: Definition of conditional proof 7 $(\vdash\psi)\Rightarrow(\vdash\phi\to\psi)$ 6: Definition of derivable rule
More examples can be found in Subsections 3.3 and 3.4.
3.2 Logic IRL
[…]
Suppose there is a countable (finite or countably infinite) set $K$ of atomic or primitive propositions. Let $A,B,C,...$ be primitive propositional variables ranging over $K$.
[…]
The propositional logic IRL is defined by the following axioms and primitive rules of inference, where $\phi,\psi,\chi,...$ are arbitrary formulas for which one can substitute any schemas.
Axioms
  1. 1.
    $\phi\leftrightarrow\phi$ (Reflexivity of bi-implication). Symmetry and transitivity can be derived.
  2. 2.
    $\phi\wedge\psi\leftrightarrow\psi\wedge\phi$, $\phi\vee\psi\leftrightarrow\psi\vee\phi$ (Commutativity).
  3. 3.
    $(\phi\wedge\psi)\wedge\chi\leftrightarrow\phi\wedge(\psi\wedge\chi)$, $(\phi\vee\psi)\vee\chi\leftrightarrow\phi\vee(\psi\vee\chi)$ (Associativity). Redundant.
  4. 4.
    $\phi\wedge(\psi\vee\chi)\leftrightarrow(\phi\wedge\psi)\vee(\phi\wedge\chi)$, $\phi\vee(\psi\wedge\chi)\leftrightarrow(\phi\vee\psi)\wedge(\phi\vee\chi)$ (Distributivity).
  5. 5.
    $\phi\wedge\top\leftrightarrow\phi$, $\phi\vee\bot\leftrightarrow\phi$ (Identity elements, top and bottom).
  6. 6.
    $\phi\wedge\neg\phi\leftrightarrow\bot$, $\phi\vee\neg\phi\leftrightarrow\top$ (Complement/Negation, LNC and LEM).
  7. 7.
    $(\phi\rightarrow\psi)\leftrightarrow(\phi\wedge\psi\leftrightarrow\phi)$ (Definition of implication).
Primitive rules of inference
  1. 1.
    $\{\phi,\psi\}\Rightarrow\phi\wedge\psi$ (Conjunction introduction).
  2. 2.
    $\phi\wedge\psi\Rightarrow\phi$ (Conjunction elimination).
  3. 3.
    $\{\phi\leftrightarrow\psi,\chi\}\Rightarrow\chi[\phi\mapsto\psi]$ (Replacement property of bi-implication).
  4. 4.
    $(\phi\vdash\psi)\Rightarrow(\vdash\phi\to\phi)$ (Rule of conditional proof). Meta Rule.
Remark 3.2.1.
  • Symmetry and transitivity of bi-implication are proved in the next subsection.
  • Associativity can be proved following a procedure similar to that by Huntington (1904).
  • In a proof, by definition, any member of its proof sequence must be either a theorem (or a theorem indicated in an outer layer), or a result from a subset of the premises and previous members in its own layer by a rule (or by a rule indicated in an outer layer), no other cases are allowed. It is important to pay attention to this, especially for a nested conditional proof (cf. Subsection 3.1).
  • Rule of conditional proof of the general form $(\Gamma\cup\{\phi\}\vdash\psi)\Rightarrow(\Gamma\vdash\phi\to\psi)$, where $\Gamma$ is a set of formulas, does not hold for IRL: if it holds, then from $\{\phi,\psi\}\vdash\phi$ it follows $\phi\vdash\psi\to\phi$, hence $\vdash\phi\to(\psi\to\phi)$, contrary to that $\nvdash_{IRL}\phi\to(\psi\to\phi)$ (cf. Subsection 5.3).
  • The logic IRL is consistent as a sub-system of classical (propositional) logic (cf. Remark 5.2.2 and Subsection 5.3).
3.3 Elementary derived results
Symmetry and transitivity of bi-implication
Rule 3.3.1 (Symmetry of bi-implication).
$\phi\leftrightarrow\psi\Rightarrow\psi\leftrightarrow\phi.$
Proof.
1     $\phi\leftrightarrow\psi$ Premise
2     $\phi\leftrightarrow\phi$ Reflexivity
3     $(\phi\leftrightarrow\phi)[\phi\mapsto\psi]$ 1,2: Replacement
4     $\phi[\phi\mapsto\psi]\leftrightarrow\phi$ 3: Selected replacement
5     $\psi\leftrightarrow\phi$ 4: Result of the replacement
6 $\phi\leftrightarrow\psi\vdash\psi\leftrightarrow\phi$ 1-5: Definition of conditional proof
7 $\phi\leftrightarrow\psi\Rightarrow\psi\leftrightarrow\phi$ 6: Definition of derivable rule
Note: The definitions of conditional proof and derivable rule can be used tacitly in a proof as usual. For example, the last two lines in the above proof can be omitted.
Rule 3.3.2 (Transitivity of bi-implication).
$\{\phi\leftrightarrow\psi,\psi\leftrightarrow\chi\}\Rightarrow\phi\leftrightarrow\chi.$
Proof.
1 $\phi\leftrightarrow\psi$ Premise
2 $\psi\leftrightarrow\chi$ Premise
3 $(\phi\leftrightarrow\psi)[\psi\mapsto\chi]$ 2,1: Replacement
4 $\phi\leftrightarrow\chi$ 3: Result of the replacement
Equivalence of bi-implication theorem and bi-inference rule
Meta Rule 3.3.1 (Bi-implication theorem to bi-inference rule).
$(\vdash\phi\leftrightarrow\psi)\Rightarrow(\phi\Leftrightarrow\psi).$
Proof.
1 $\vdash\phi\leftrightarrow\psi$ Premise (begin meta proof)
2     $\phi$ Premise (begin proof 1)
3     $\phi\leftrightarrow\psi$ 1 (outside assumed theorem)
4     $\phi[\phi\mapsto\psi]$ 3,2: Replacement
5     $\psi$ 4: Result of the replacement (end proof 1)
6 $\phi\vdash\psi$ 2–5: Definition of conditional proof
7 $\phi\Rightarrow\psi$ 6: Definition of derivable rule
8     $\psi$ Premise (begin proof 2)
9     $\phi\leftrightarrow\psi$ 1 (outside assumed theorem)
10     $\psi\leftrightarrow\phi$ 9: Symmetry of bi-implication
11     $\psi[\psi\mapsto\phi]$ 10,8: Replacement
12     $\phi$ 11: Result of the replacement (end proof 2)
13 $\psi\vdash\phi$ 8–12: Definition of conditional proof
14 $\psi\Rightarrow\phi$ 13: Definition of derivable rule
15 $\psi\Leftrightarrow\phi$ 7,14: Definition of bi-inference rule (end meta proof)
By this meta rule, from bi-implication axioms we obtain bi-inference rules directly. For example, from $\phi\wedge\psi\leftrightarrow\psi\wedge\phi$, $(\phi\wedge\psi)\wedge\chi\leftrightarrow\phi\wedge(\psi\wedge\chi)$ and $(\phi\to\psi)\leftrightarrow(\phi\wedge\psi\leftrightarrow\phi)$, we obtain rules $\phi\wedge\psi\Leftrightarrow\psi\wedge\phi$, $(\phi\wedge\psi)\wedge\chi\Leftrightarrow\phi\wedge(\psi\wedge\chi)$ and $\phi\to\psi\Leftrightarrow\phi\wedge\psi\leftrightarrow\phi$, respectively.
Rule 3.3.3 (Bi-implication introduction).
$\{\phi\to\psi,\psi\to\phi\}\Rightarrow\phi\leftrightarrow\psi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\psi\to\phi$ Premise
3 $\phi\wedge\psi\leftrightarrow\phi$ 1: Definition of $\to$
4 $\phi\leftrightarrow\phi\wedge\psi$ 3: Symmetry of $\leftrightarrow$
5 $\psi\wedge\phi\leftrightarrow\psi$ 2: Definition of $\to$
6 $\phi\wedge\psi\leftrightarrow\psi\wedge\phi$ Commutativity
7 $\phi\wedge\psi\leftrightarrow\psi$ 6,5: Transitivity of $\leftrightarrow$
8 $\phi\leftrightarrow\psi$ 4,7: Transitivity of $\leftrightarrow$
Meta Rule 3.3.2 (Bi-inference rule to bi-implication theorem).
$(\phi\Leftrightarrow\psi)\Rightarrow(\vdash\phi\leftrightarrow\psi).$
Proof.
1 $\phi\Leftrightarrow\psi$ Premise
2       $\phi$ Premise
3       $\psi$ 2: The rule assumed in 1
4     $\phi\to\psi$ 2–3: Conditional proof
5     $\psi\to\phi$ 1–4: Similarly
6     $\phi\leftrightarrow\psi$ 4,5: Bi-implication introduction
7 $\vdash\phi\leftrightarrow\psi$ 4–6: Unconditional proof
Corollary 3.3.1 (Equivalence of bi-implication theorem and bi-inference rule).
$(\vdash\phi\leftrightarrow\psi)\Leftrightarrow(\phi\Leftrightarrow\psi).$
Proof.
It follows from Meta Rules 3.3.1 and 3.3.2 by the definition of bi-inference rule. ∎
Thus, we do not need to distinguish bi-implication theorem $\vdash\phi\leftrightarrow\psi$ and bi-inference rule $\phi\Leftrightarrow\psi$: if we have one of the two, we have both.
Tacit use of symmetry properties
Meta Rule 3.3.3.
Let $F(\phi\circ\phi)$ be a schema containing $\phi\circ\psi$, where $\circ$ stands for $\leftrightarrow$, $\wedge$, or $\vee$. Then
$F(\phi\circ\psi)\Leftrightarrow F(\psi\circ\phi).$
Proof.
We already have $(\phi\circ\psi)\leftrightarrow(\psi\circ\phi)$ for $\leftrightarrow$, $\wedge$ and $\vee$, so the proof can be done below:
1     $F(\phi\circ\psi)$ Premise 2     $(\phi\circ\psi)\leftrightarrow(\psi\circ\phi)$ Known theorem 3     $F(\psi\circ\phi)$ 2,1: Replacement 4 $F(\phi\circ\psi)\Rightarrow F(\psi\circ\phi)$ 1–3: Definition of derivable rule 5 $F(\psi\circ\phi)\Rightarrow F(\phi\circ\psi)$ 1–4: Similarly 6 $F(\psi\circ\phi)\Leftrightarrow F(\phi\circ\psi)$ 4,5: Definition of bi-inference rule.
Thus, the symmetry properties of $\leftrightarrow$, $\wedge$ and $\vee$ can be used tacitly in a proof, viewing $F(\phi\circ\psi)$ and $F(\psi\circ\phi)$ as the same thing.
Modus ponens
Rule 3.3.4 (Modus ponens).
$\{\phi\to\psi,\phi\}\Rightarrow\psi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\phi$ Premise
3 $\phi\leftrightarrow\phi\wedge\psi$ 1: Definition of implication
4 $\phi\wedge\psi$ 3,2: Replacement
5 $\psi$ 4: Conjunction elimination
Equivalence of the three forms
With the rules of commutativity and associativity for conjunction, $\phi_{1}\wedge\cdots\wedge\phi_{n}$ is well defined for any $n\geq 1$. And from the bidirectional reflexivity rule $\phi\Leftrightarrow\phi$ we get its unidirectional form $\phi\Rightarrow\phi$ by the definition of bi-inference rule. Therefore, the rules of conjunction introduction and elimination can be generalized to $\{\phi_{1},...,\phi_{n}\}\Rightarrow\phi_{1}\wedge\cdots\wedge\phi_{n}$ and $\phi_{1}\wedge\cdots\wedge\phi_{n}\Rightarrow\phi_{i}$ respectively, for all finite $n\geq 1$.
Meta Rule 3.3.4 (Rule of conditional proof with multiple premises).
($(\{\phi_{1},...,\phi_{n}\}\vdash\psi)\Rightarrow(\vdash\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi).$
Proof.
Suppose $\{\phi_{1},...,\phi_{n}\}\vdash\psi$ has a proof sequence $(\psi_{1},...,\psi)$, then $(\phi_{1},...,\phi_{n},\psi_{1},...,\psi)$ is a proof sequence of $\phi_{1}\wedge\cdots\wedge\phi_{n}\vdash\psi$ as shown below:
0 $\phi_{1}\wedge\cdots\wedge\phi_{n}$ Premise 1 $\phi_{1}$ 0: Conjunction elimination $\vdots$ n $\phi_{n}$ 0: Conjunction elimination $\psi_{1}$ 1…n: First proof-sequence member of $\{\phi_{1},...,\phi_{n}\}\vdash\psi$vdots $\psi$ 1…n: Last proof-sequence member of $\{\phi_{1},...,\phi_{n}\}\vdash\psi$
Thus, from $\phi_{1}\wedge\cdots\wedge\phi_{n}\vdash\psi$ we obtain $\vdash\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi$ by the rule of conditional proof with single premise. ∎
Meta Rule 3.3.5.
(Converse of CP rule)
$(\vdash\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi)\Rightarrow(\{\phi_{1},...,\phi_{n}\}\vdash\psi).$
Proof.
1 $\vdash\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi$ Premise (meta proof)
2     $\phi_{1}$ Premise (proof)
    $\vdots$
3     $\phi_{n}$ Premise (proof)
4     $\phi_{1}\wedge\cdots\wedge\phi_{n}$ 2…3: Conjunction introduction
5     $\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi$ 1 (assumed theorem)
6     $\psi$ 5,4: Modus ponens
7 $\{\phi_{1},...,\phi_{n}\}\vdash\psi$ 2–6: Definition of conditional proof
Meta Rule 3.3.6 (Equivalence of the three forms).
$(\vdash\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi)\Leftrightarrow(\{\phi_{1},...,\phi_{n}\}\vdash\psi)\Leftrightarrow(\{\phi_{1},...,\phi_{n}\}\Rightarrow\psi).$
Proof.
It follows from the definition of derivable rule, Meta Rule 3.3.4 and Meta Rule 3.3.5, and the definition of bi-inference rule. ∎
Therefore, if we have any one of the three forms, we have the other two. So we just need to find either an unconditional proof $\vdash\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi$ or a conditional proof $\{\phi_{1},...,\phi_{n}\}\vdash\psi$ to obtain both the theorem $\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi$ and the rule $\{\phi_{1},...,\phi_{n}\}\Rightarrow\psi$.
Other elementary results
Rule 3.3.5 (Leibniz’s replacement).
$\phi\leftrightarrow\psi\Rightarrow\chi\leftrightarrow\chi[\phi\mapsto\psi].$
Proof.
1 $\phi\leftrightarrow\psi$ Premise
2 $\chi\leftrightarrow\chi$ Reflexivity
3 $(\chi\leftrightarrow\chi)[\phi\mapsto\psi]$ 1,2: Replacement
4 $\chi\leftrightarrow\chi[\phi\mapsto\psi]$ 3: Selected replacement
Theorem 3.3.1 (Top is a theorem).
$\top$.⊤ .
Proof.
1 $\phi\leftrightarrow\phi$ Reflexivity
2 $(\phi\leftrightarrow\phi)\wedge\top$ 1: Identity element
3 $\top$ 2: Conjunction elimination
Meta Rule 3.3.7 (Theorem like top).
$(\vdash\psi)\Rightarrow(\vdash\phi\to\psi).$
Proof.
1 $\vdash\psi$ Premise (meta proof)
2     $\phi$ Premise (proof)
3     $\psi$ 1 (assumed theorem)
4 $\vdash\phi\to\psi$ 2–3: Conditional proof
A theorem functions like the top $\top$.
Meta Rule 3.3.8 (Theorem equivalence).
$\{\vdash\phi,\vdash\psi\}\Rightarrow(\vdash\phi\leftrightarrow\psi).$
Proof.
1 $\vdash\phi$ Premise
2 $\vdash\psi$ Premise
3 $\vdash\psi\to\phi$ 1: Theorem like top
4 $\vdash\phi\to\psi$ 2: Theorem like top
5     $\psi\to\phi$ 3 (theorem from an outer layer)
6     $\phi\to\psi$ 4 (theorem from an outer layer)
7     $\phi\leftrightarrow\psi$ 5,6: Bi-implication introduction
8 $\vdash\phi\leftrightarrow\psi$ 5–7: Unconditional proof
Meta Rule 3.3.9 (Theorem representation).
$(\vdash\phi\leftrightarrow\top)\Leftrightarrow(\vdash\phi).$
Proof.
1 $\vdash\phi\leftrightarrow\top$ Premise
2     ϕitalic-$\phi\leftrightarrow\top$ 1 (assumed theorem)
3     $\top$ Top is a theorem
4     $\phi$ 2,3: Replacement $[\top\mapsto\phi]$
5 $\vdash\phi$ 2–4: Unconditional proof
1 $\vdash\phi$ Premise
2     $\phi$ 1 (assumed theorem)
3     $\top$ Top is a theorem
4     ϕitalic-$\phi\leftrightarrow\top$ 2,3: Theorem equivalence
5 $\vdash\phi\leftrightarrow\top$ 2–4: Unconditional proof
For example, $\phi\vee\neg\phi\leftrightarrow\top$ and $\phi\vee\neg\phi$ are equivalent representations of LEM as a theorem.
3.4 More derivations in IRL
Some basic theorems or rules
Note that no matter what is proved is a theorem or a rule, we have both a theorem and a rule according to the equivalence of the three forms (Meta Rule 3.3.6) and the equivalence of bi-implication theorem and bi-inference rule (Corollary 3.3.1).
Theorem 3.4.1 (Complementarity of top and bottom).
[…]
Proof.
[…]
The second is proved dually. ∎
Theorem 3.4.2 (Double negation).
$\neg\neg\phi\leftrightarrow\phi.$
Proof.
$\neg\neg\phi$
1     $\leftrightarrow\neg\neg\phi\wedge\top$ Identity element
2     $\leftrightarrow\neg\neg\phi\wedge(\phi\vee\neg\phi)$ LEM: Leibniz’s replacement
3     ($\leftrightarrow(\neg\neg\phi\wedge\phi)\vee(\neg\neg\phi\wedge\neg\phi)$ Distributivity
4     ($\leftrightarrow(\neg\neg\phi\wedge\phi)\vee\bot$ LNC: Leibniz’s replacement
5     ($\leftrightarrow(\neg\neg\phi\wedge\phi)\vee(\neg\phi\wedge\phi)$ LNC: Leibniz’s replacement
6     ($\leftrightarrow(\neg\neg\phi\vee\neg\phi)\wedge\phi$ Distributivity
7     $\leftrightarrow\top\wedge\phi$ LEM: Leibniz’s replacement
8     $\leftrightarrow\phi$ Identity element
9 $\neg\neg\phi\leftrightarrow\phi$ 1–8: Transitivity of bi-implication
Note: Leibniz’s replacement may be used tacitly, e.g. “LEM: Leibniz’s replacement” can be written simply as “LEM” in a reason of a proof line.
Theorem 3.4.3 (Idempotence).
$\phi\wedge\phi\leftrightarrow\phi,\;\phi\vee\phi\leftrightarrow\phi.$
Proof.
$\phi\wedge\phi$
1     $\leftrightarrow(\phi\wedge\phi)\vee\bot$ Identity
2     $\leftrightarrow(\phi\wedge\phi)\vee(\phi\wedge\neg\phi)$ LNC
3     $\leftrightarrow\phi\wedge(\phi\vee\neg\phi)$ Distributivity
4     $\leftrightarrow\phi\wedge\top$ LEM
5     $\leftrightarrow\phi$ Identity
6 $\phi\wedge\phi\leftrightarrow\phi$ 1–5: Transitivity
The second is proved similarly. ∎
Theorem 3.4.4 (Annihilator).
$\bot\wedge\phi\leftrightarrow\bot,\;\top\vee\phi\leftrightarrow\top.$
Proof.
$\bot\wedge\phi$
1     $\leftrightarrow(\bot\wedge\phi)\vee\bot$ Identity
2     $\leftrightarrow(\bot\wedge\phi)\vee(\neg\phi\wedge\phi)$ LNC
3     ($\leftrightarrow(\bot\vee\neg\phi)\wedge\phi$ Distributivity
4     $\leftrightarrow\neg\phi\wedge\phi$ Identity
[…]
6 $\bot\wedge\phi\leftrightarrow\bot$ 1–5: Transitivity
The second is proved similarly. ∎
Theorem 3.4.5 (Absorption laws).
$\phi\wedge(\phi\vee\psi)\leftrightarrow\phi,\;\phi\vee(\phi\wedge\psi)\leftrightarrow\phi.$
Proof.
$\phi\wedge(\phi\vee\psi)$
1     $\leftrightarrow(\phi\vee\bot)\wedge(\phi\vee\psi)$ Identity
2     $\leftrightarrow\phi\vee(\bot\wedge\psi)$ Distributivity
3     $\leftrightarrow\phi\vee\bot$ Annihilator
4     $\leftrightarrow\phi$ Identity
5 $\phi\wedge(\phi\vee\psi)\leftrightarrow\phi$ 1–4: Transitivity
The second is proved similarly. ∎
Basic properties of implication
Theorem 3.4.6 (Reflexivity of implication).
$\phi\to\phi.$
Proof.
1 $\phi\wedge\phi\leftrightarrow\phi$ Idempotence
2 $\phi\to\phi$ 1: Definition of implication
Antisymmetry of implication is the same as Bi-implication introduction (Rule 3.3.3).
Rule 3.4.1 (Transitivity of implication).
$\{\phi\to\psi,\psi\to\chi\}\Rightarrow\phi\to\chi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\psi\to\chi$ Premise
3 $\phi\wedge\psi\leftrightarrow\phi$ 1: Definition of implication
4 $\psi\wedge\chi\leftrightarrow\psi$ 2: Definition of implication
$\phi\wedge\chi$
5     $\leftrightarrow\phi\wedge\psi\wedge\chi$ 3: Replacement $[\phi\mapsto\phi\wedge\psi]$
6     $\leftrightarrow\phi\wedge\psi$ 4: Replacement $[\psi\wedge\chi\mapsto\psi]$
7     $\leftrightarrow\phi$ 3: Replacement
8 $\phi\wedge\chi\leftrightarrow\phi$ 5–7: Transitivity of bi-implication
9 $\phi\to\chi$ 8: Definition of implication
Equivalent definitions of implication
Lemma 3.4.1.
$\phi\wedge(\neg\phi\vee\psi)\leftrightarrow\phi\wedge\psi,\;\phi\vee(\neg\phi\wedge\psi)\leftrightarrow\phi\vee\psi.$
Proof.
$\phi\wedge(\neg\phi\vee\psi)$
1     $\leftrightarrow(\phi\wedge\neg\phi)\vee(\phi\wedge\psi)$ Distributivity
2     $\leftrightarrow\bot\vee(\phi\wedge\psi)$ LNC
3     $\leftrightarrow\phi\wedge\psi$ Identity
4 $\phi\wedge(\neg\phi\vee\psi)\leftrightarrow\phi\wedge\psi$ 1–3: Transitivity
The second is proved dually. ∎
Rule 3.4.2 (Equivalent definitions of implication).
$\phi\to\psi$
$\Leftrightarrow$ $\phi\wedge\psi\leftrightarrow\phi$
$\Leftrightarrow$ $\phi\vee\psi\leftrightarrow\psi$
$\Leftrightarrow$ $\phi\wedge\neg\psi\leftrightarrow\bot$
$\Leftrightarrow$ $\neg\phi\vee\psi\leftrightarrow\top.$
Proof.
1 $\phi\to\psi$ Premise
2 $\phi\wedge\psi\leftrightarrow\phi$ 1: Definition of implication
1 $\phi\wedge\psi\leftrightarrow\phi$ Premise
$\phi\vee\psi$
2     $\leftrightarrow(\phi\wedge\psi)\vee\psi$ 1: Replacement $[\phi\mapsto\phi\wedge\psi]$
3     $\leftrightarrow\psi$ Absorption
4 $\phi\vee\psi\leftrightarrow\psi$ 2–3: Transitivity
1 $\phi\vee\psi\leftrightarrow\psi$ Premise
$\phi\wedge\neg\psi$
2     $\leftrightarrow(\phi\wedge\neg\psi)\vee\bot$ Identity
3     $\leftrightarrow(\phi\wedge\neg\psi)\vee(\psi\wedge\neg\psi)$ LNC
4     $\leftrightarrow(\phi\vee\psi)\wedge\neg\psi$ Distributivity
5     $\leftrightarrow\psi\wedge\neg\psi$ 1: Replacement $[\phi\vee\psi\mapsto\psi]$
[…]
7 $\phi\wedge\neg\psi\leftrightarrow\bot$ 2–6: Transitivity
1 $\phi\wedge\neg\psi\leftrightarrow\bot$ Premise
$\neg\phi\vee\psi$
2     $\leftrightarrow\neg\phi\vee\bot\vee\psi$ Identity
3     $\leftrightarrow\neg\phi\vee(\phi\wedge\neg\psi)\vee\psi$ 1: Replacement $[\bot\mapsto\phi\wedge\neg\psi]$
4     $\leftrightarrow\neg\phi\vee\phi\vee\psi$ Lemma 3.4.1
5     $\leftrightarrow\top\vee\psi$ LEM
[…]
7 $\neg\phi\vee\psi\leftrightarrow\top$ 2–6: Transitivity
1 $\neg\phi\vee\psi\leftrightarrow\top$ Premise
$\phi\wedge\psi$
2     $\leftrightarrow\bot\vee(\phi\wedge\psi)$ Identity
3     $\leftrightarrow(\phi\wedge\neg\phi)\vee(\phi\wedge\psi)$ LNC
4     $\leftrightarrow\phi\wedge(\neg\phi\vee\psi)$ Distributivity
5     $\leftrightarrow\phi\wedge\top$ 1: Replacement
6     $\leftrightarrow\phi$ Identity
7 $\phi\wedge\psi\leftrightarrow\phi$ 2–6: Transitivity
1 $\phi\wedge\psi\leftrightarrow\phi$ Premise
2 $\phi\to\psi$ 1: Definition of implication
Hereinafter, any one of the four is simply called the definition of implication (by bi-implication).
Implication-to-Disjunction
Theorem 3.4.7 (Implication-to-Disjunction / Upper bound of implication).
$(\phi\to\psi)\to\neg\phi\vee\psi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\neg\phi\vee\psi\leftrightarrow\top$ 1: Definition of implication
3 $\top$ Top is a theorem
4 $\neg\phi\vee\psi$ 2,3: Replacement [$[\top\mapsto\neg\phi\vee\psi]$
De Morgan’s laws
Theorem 3.4.8 (De Morgan’s laws).
$\neg(\phi\wedge\psi)\leftrightarrow(\neg\phi\vee\neg\psi),\;\neg(\phi\vee\psi)\leftrightarrow(\neg\phi\wedge\neg\psi).$
Proof.
$\neg\neg(\phi\wedge\psi)\vee(\neg\phi\vee\neg\psi)$
1     $\leftrightarrow(\phi\wedge\psi)\vee(\neg\phi\vee\neg\psi)$ Double negation
2     $\leftrightarrow(\phi\vee\neg\phi\vee\neg\psi)\wedge(\psi\vee\neg\phi\vee\neg\psi)$ Distributivity
3     ($\leftrightarrow(\top\vee\neg\psi)\wedge(\top\vee\neg\phi)$ LEM
[…]
6 $\neg\neg(\phi\wedge\psi)\vee(\neg\phi\vee\neg\psi)\leftrightarrow\top$ 1–5: Transitivity
7 $\neg(\phi\wedge\psi)\to(\neg\phi\vee\neg\psi)$ 6: Definition of implication
($(\neg\phi\vee\neg\psi)\wedge\neg\neg(\phi\wedge\psi)$
8     ($\leftrightarrow(\neg\phi\vee\neg\psi)\wedge(\phi\wedge\psi)$ Double negation
9     ($\leftrightarrow(\neg\phi\wedge\phi\wedge\psi)\vee(\neg\psi\wedge\phi\wedge\psi)$ Distributivity
10     $\leftrightarrow(\bot\wedge\psi)\vee(\bot\wedge\phi)$ LNC
[…]
13 ($(\neg\phi\vee\neg\psi)\wedge\neg\neg(\phi\wedge\psi)\leftrightarrow\bot$ 8–12: Transitivity
14 $\neg\phi\vee\neg\psi\to\neg(\phi\wedge\psi)$ 13: Definition of implication
15 $\neg(\phi\wedge\psi)\leftrightarrow(\neg\phi\vee\neg\psi)$ 7,14: Bi-implication introduction
The second is proved dually. ∎
Lattice-like properties
Theorem 3.4.9 (Conjunction elimination and disjunction introduction).
$\phi\wedge\psi\to\phi,\;\phi\to\phi\vee\psi.$
Proof.
In terms of equivalent definitions of implication, conjunction elimination and disjunction introduction are equivalent to absorption laws $\phi\vee(\phi\wedge\psi)\leftrightarrow\phi$ and $\phi\wedge(\phi\vee\psi)\leftrightarrow\phi$ respectively. ∎
Rule 3.4.3 (One-to-two and Two-to-one).
$\{\chi\to\phi,\chi\to\psi\}\Leftrightarrow\chi\to\phi\wedge\psi,\;\{\phi\to\chi,\psi\to\chi\}\Leftrightarrow\phi\vee\psi\to\chi.$
Proof.
1 $\chi\to\phi$ Premise
2 $\chi\to\psi$ Premise
3 $\chi\wedge\phi\leftrightarrow\chi$ 1: Definition of implication
4 $\chi\wedge\psi\leftrightarrow\chi$ 2: Definition of implication
5 $\chi\wedge\phi\wedge\psi\leftrightarrow\chi$ 3,4: Replacement $[\chi\mapsto\chi\wedge\phi]$
6 $\chi\to\phi\wedge\psi$ 5: Definition of implication
1 $\chi\to\phi\wedge\psi$ Premise
2 $\phi\wedge\psi\to\phi$ Conjunction elimination
3 $\chi\to\phi$ 1–2: Transitivity
4 $\chi\to\psi$ 1–3: Similarly
The second is proved dually. ∎
Rule 3.4.4 (Strengthening).
$\phi\to\psi\Rightarrow\phi\to\phi\wedge\psi,\;\phi\to\psi\Rightarrow\phi\vee\psi\to\psi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\phi\to\phi$ Reflexivity
3 $\phi\to\phi\wedge\psi$ 1,2: One-to-two
The second is proved dually. ∎
Rule 3.4.5 (Weakening).
$\phi\to\psi\Rightarrow\phi\wedge\chi\to\psi,\;\phi\to\psi\Rightarrow\phi\to\psi\vee\chi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\phi\wedge\chi\to\phi$ Conjunction elimination
3 $\phi\wedge\chi\to\psi$ 2,1: Transitivity
The second is proved dually. ∎
Rule 3.4.6 (Compatibility of implication with conjunction and disjunction).
$\phi\to\psi\Rightarrow\phi\wedge\chi\to\psi\wedge\chi,\;\phi\to\psi\Rightarrow\phi\vee\chi\to\psi\vee\chi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\phi\wedge\chi\to\psi$ 1: Weakening
3 $\phi\wedge\chi\to\chi$ Conjunction elimination
4 $\phi\wedge\chi\to\psi\wedge\chi$ 2,3: One-to-two
The second is proved dually. ∎
Rule 3.4.7 (Monotonicity of conjunction and disjunction w.r.t. implication).
$\{\phi_{1}\to\psi_{1},\phi_{2}\to\psi_{2}\}\Rightarrow\phi_{1}\wedge\phi_{2}\to\psi_{1}\wedge\psi_{2},$
$\{\phi_{1}\to\psi_{1},\phi_{2}\to\psi_{2}\}\Rightarrow\phi_{1}\vee\phi_{2}\to\psi_{1}\vee\psi_{2}.$
Proof.
1 $\phi_{1}\to\psi_{1}$ Premise
2 $\phi_{2}\to\psi_{2}$ Premise
3 $\phi_{1}\wedge\phi_{2}\to\psi_{1}$ 1: Weakening
4 $\phi_{1}\wedge\phi_{2}\to\psi_{2}$ 2: Weakening
5 $\phi_{1}\wedge\phi_{2}\to\psi_{1}\wedge\psi_{2}$ 3,4: One-to-two
The second is proved dually. ∎
Theorem 3.4.10 (Contraposition).
$(\phi\to\psi)\leftrightarrow(\neg\psi\to\neg\phi).$
Proof.
$(\phi\to\psi)$
1     $\leftrightarrow(\phi\wedge\neg\psi\leftrightarrow\bot)$ Definition of implication
2     ($\leftrightarrow(\neg\psi\wedge\neg\neg\phi\leftrightarrow\bot)$ Commutativity, Double negation
3     ($\leftrightarrow(\neg\psi\to\neg\phi)$ Definition of implication
4 $(\phi\to\psi)\leftrightarrow(\neg\psi\to\neg\phi)$ 1–3: Transitivity
Rule 3.4.8 (Modus tollens).
$\{\phi\to\psi,\neg\psi\}\Rightarrow\neg\phi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\neg\psi$ Premise
3 $\neg\psi\to\neg\phi$ 1: Contraposition
4 $\neg\phi$ 2,3: Modus ponens
Theorem 3.4.11 (ECQ and its dual).
$\bot\to\phi,\;\phi\to\top.$
Proof.
In terms of equivalent definitions of implication, ECQ and its dual are equivalent to Annihilators $\bot\wedge\phi\leftrightarrow\bot$ and $\top\vee\phi\leftrightarrow\top$ respectively. ∎
Some properties of bi-implication
Rule 3.4.9 (Bi-implication elimination).
$\phi\leftrightarrow\psi\Rightarrow\phi\to\psi.$
Proof.
1 $\phi\leftrightarrow\psi$ Premise
2 $\phi\to\phi$ Reflexivity of implications
3 $\phi\to\psi$ 1,2: Replacement
Rule 3.4.10 (Bi-implication by implications).
$\phi\leftrightarrow\psi\Leftrightarrow(\phi\to\psi)\wedge(\psi\to\phi).$
Proof.
1 $\phi\leftrightarrow\psi$ Premise
2 $\phi\to\psi$ 1: Bi-implication elimination
3 $\psi\to\phi$ 1: Bi-implication elimination
4 $(\phi\to\psi)\wedge(\psi\to\phi)$ 2,3: Conjunction introduction
1 $(\phi\to\psi)\wedge(\psi\to\phi)$ Premise
2 $\phi\leftrightarrow\psi$ 1: Bi-implication introduction
Rule 3.4.11 (Bi-implication of negation).
$\phi\leftrightarrow\psi\Leftrightarrow\neg\phi\leftrightarrow\neg\psi.$
Proof.
$(\phi\leftrightarrow\psi)$
1     $\leftrightarrow(\phi\to\psi)\wedge(\psi\to\phi)$ Bi-implication by implications
2     ($\leftrightarrow(\neg\psi\to\neg\phi)\wedge(\neg\phi\to\neg\psi)$ Contraposition
3     ($\leftrightarrow(\neg\phi\leftrightarrow\neg\psi)$ Bi-implication by implications
More properties of implication
Theorem 3.4.12 (Generalized ECQ of implication).
$(\phi\to\psi)\to(\chi\to(\phi\to\psi)).$
Proof.
1 $\phi\to\psi$ Premise
2 $\phi\wedge\psi\leftrightarrow\phi$ 1: Definition of implication
3 $\phi\wedge\psi\to\psi$ Conjunction elimination
4 $\chi\to(\phi\wedge\psi\to\psi)$ 3: Theorem like top
5 $\chi\to(\phi\to\psi)$ 2,4: Replacement $[\phi\wedge\psi\mapsto\phi]$
It can also be derived in a meta proof:
1 $\vdash\phi\wedge\psi\to\psi$ Conjunction elimination 2 $\vdash\chi\to(\phi\wedge\psi\to\psi)$ 1: Theorem like top 3     $\phi\to\psi$ Premise 4     $\phi\wedge\psi\leftrightarrow\phi$ 3: Definition of implication 5     $\chi\to(\phi\wedge\psi\to\psi)$ 2 (theorem from an outer layer) 6     $\chi\to(\phi\to\psi)$ 4,5: Replacement $[\phi\wedge\psi\mapsto\phi]$ 7 $\vdash(\phi\to\psi)\to(\chi\to(\phi\to\psi))$ 3–6: Conditional proof
Note that Generalized ECQ of the general form $\psi\to(\phi\to\psi)$ does not hold for IRL (cf. Subsection 5.3).
Theorem 3.4.13 (Safe generalized ECQ).
$(\top\to\phi)\to(\psi\to\phi),\;(\phi\to\bot)\to(\phi\to\psi).$
Proof.
1 $\top\to\phi$ Premise
2 $\psi\to\top$ Dual of ECQ
3 $\psi\to\phi$ 2,1: Transitivity
The second is proved dually.
Theorem 3.4.14 (Importation).
$(\phi\to(\psi\to\chi))\to(\phi\wedge\psi\to\chi).$
Proof.
1 $\phi\to(\psi\to\chi)$ Premise
$\phi\wedge\psi$
2     $\to(\psi\to\chi)\wedge\psi$ 1: Compatibility of $\to$ with $\wedge$
3     $\to\chi$ Modus ponens
4 $\phi\wedge\psi\to\chi$ 2–3: Transitivity
Corollary 3.4.1 (Contraction).
$(\phi\to(\phi\to\psi))\to(\phi\to\psi)$.
Proof.
It follows from the above theorem and Idempotence. ∎
Exportation (converse of importation) $(\phi\wedge\psi\to\chi)\to(\phi\to(\psi\to\chi))$ does not hold for IRL except for some special cases as follows.
Theorem 3.4.15 (Exported antisymmetry of implication / bi-implication introduction).
$(\phi\to\psi)\to((\psi\to\phi)\to(\phi\leftrightarrow\psi)).$
Proof.
1 $\phi\to\psi$ Premise
2 $(\psi\to\phi)\to(\phi\to\psi)$ 1: Generalized ECQ of implication
3 $(\psi\to\phi)\to(\phi\to\psi)\wedge(\psi\to\phi)$ 2: Strengthening
4 $(\phi\to\psi)\wedge(\psi\to\phi)\to(\phi\leftrightarrow\psi)$ Bi-implication introduction
5 $(\psi\to\phi)\to(\phi\leftrightarrow\psi)$ 3,4: Transitivity
Theorem 3.4.16 (Exported transitivity of implication / Suffixing and Prefixing).
$(\phi\to\psi)\to((\psi\to\chi)\to(\phi\to\chi)),\;(\phi\to\psi)\to((\chi\to\phi)\to(\chi\to\psi)).$
Proof.
1 $\phi\to\psi$ Premise
2 $(\psi\to\chi)\to(\phi\to\psi)$ 1: Generalized ECQ of implication
3 $(\psi\to\chi)\to(\phi\to\psi)\wedge(\psi\to\chi)$ 2: Strengthening
4 $(\phi\to\psi)\wedge(\psi\to\chi)\to(\phi\to\chi)$ Transitivity of implication
5 $(\psi\to\chi)\to(\phi\to\chi)$ 3–4: Transitivity of implication
The second is obtained via uniform substitution and contraposition.
Theorem 3.4.17 (Exported one-to-two and two-to-one).
$(\chi\to\phi)\to((\chi\to\psi)\to(\chi\to\phi\wedge\psi)),\;(\phi\to\chi)\to((\psi\to\chi)\to(\phi\vee\psi\to\chi)).$
Proof.
1 $\chi\to\phi$ Premise
2 $(\chi\to\psi)\to(\chi\to\phi)$ 1: Generalized ECQ of implication
3 $(\chi\to\psi)\to(\chi\to\phi)\wedge(\chi\to\psi)$ 2: Strengthening
4 $(\chi\to\phi)\wedge(\chi\to\psi)\to(\chi\to\phi\wedge\psi)$ One-to-two
5 $(\chi\to\psi)\to(\chi\to\phi\wedge\psi))$ 3–4: Transitivity
The second is obtained via contraposition.
Theorem 3.4.18 (Self-distributivity of implication).
$(\chi\to(\phi\to\psi))\to((\chi\to\phi)\to(\chi\to\psi)).$
Proof.
1 $\chi\to(\phi\to\psi)$ Premise
2 $\chi\wedge\phi\to\psi$ 1: Importation
3 $(\chi\to\phi)\to(\chi\wedge\phi\to\psi)$ 2: Generalized ECQ of implication
4 $(\chi\to\phi)\to(\chi\to\chi\wedge\phi)$ Strengthening
5 $(\chi\to\phi)\to(\chi\to\chi\wedge\phi)\wedge(\chi\wedge\phi\to\psi)$ 4,3: One-to-two
6 $(\chi\to\chi\wedge\phi)\wedge(\chi\wedge\phi\to\psi)\to(\chi\to\psi)$ Transitivity
7 $(\chi\to\phi)\to(\chi\to\psi)$ 5,6: Transitivity
Implication with one or no variable
Theorem 3.4.19 (Univariate implication expressions).
$(\phi\to\neg\phi)\leftrightarrow(\phi\leftrightarrow\bot)\leftrightarrow(\phi\to\bot);$
($(\neg\phi\to\phi)\leftrightarrow(\phi\leftrightarrow\top)\leftrightarrow(\top\to\phi).$
Proof.
$(\phi\to\neg\phi)$
1     $\leftrightarrow(\phi\wedge\neg\neg\phi\leftrightarrow\bot)$ Definition of implication
2     $\leftrightarrow(\phi\wedge\phi\leftrightarrow\bot)$ Double negation
3     $\leftrightarrow(\phi\leftrightarrow\bot)$ Idempotence
4     $\leftrightarrow(\phi\to\bot)\wedge(\bot\to\phi)$ Bi-implication by implication
5     (ϕ)absentitalic-$\leftrightarrow(\phi\to\bot)$ Theorem (ECQ) like top
6 $(\phi\to\neg\phi)\leftrightarrow(\phi\leftrightarrow\bot)\leftrightarrow(\phi\to\bot)$ 1–3–5: Transitivity
The second is proved dually. ∎
Theorem 3.4.20 (Upper bound of univariate implication).
$(\top\to\phi)\to\phi,\;(\phi\to\bot)\to\neg\phi.$
Proof.
They follow from the upper bound of implication $(\phi\to\psi)\to\neg\phi\vee\psi$ (Theorem 3.4.7) by uniform substitution, and properties of top and bottom. ∎
Corollary 3.4.2 (Contradictory implication).
[…]
Proof.
It follows from Theorem 3.4.20 by uniform substitution. ∎
Proof methods related
Rule 3.4.12 (Constructive dilemma).
$\{\phi_{1}\vee\phi_{2},\phi_{1}\to\psi_{1},\phi_{2}\to\psi_{2}\}\Rightarrow\psi_{1}\vee\psi_{2}.$
Proof.
1 $\phi_{1}\vee\phi_{2}$ Premise
2 $\phi_{1}\to\psi_{1}$ Premise
3 $\phi_{2}\to\psi_{2}$ Premise
4 $\phi_{1}\vee\phi_{2}\to\psi_{1}\vee\psi_{2}$ 2,3: Monotonicity of $\vee$ w.r.t. $\to$
5 $\psi_{1}\vee\psi_{2}$ 1,4: Modus penens
Rule 3.4.13 (Proof by exhaustion).
$\{\phi\vee\psi,\phi\to\chi,\psi\to\chi\}\Rightarrow\chi.$
Proof.
1 $\phi\vee\psi$ Premise
2 $\phi\to\chi$ Premise
3 $\psi\to\chi$ Premise
4 $\phi\vee\psi\to\chi$ 2,3: Two-to-one
5 $\chi$ 1,4: Modus ponens
Rule 3.4.14 (LEM exhaustion).
$\{\phi\to\psi,\neg\phi\to\psi\}\Rightarrow\psi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\neg\phi\to\psi$ Premise
3 $\psi\vee\neg\phi$ LEM
4 $\psi$ 3,1,2: Proof by exhaustion
Rule 3.4.15 (Reductio ad absurdum (RAA)).
$\{\phi\to\psi,\phi\to\neg\psi\}\to\neg\phi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\phi\to\neg\psi$ Premise
3 $\neg\psi\to\neg\phi$ 1: Contraposition
4 $\psi\to\neg\phi$ 2: Contraposition, Double negation
5 $\neg\phi$ 3,4: LEM exhaustion
The rules $\{\phi\to\psi,\phi\to\neg\psi\}\Rightarrow\neg\phi$, $\phi\to\neg\phi\Rightarrow\neg\phi$ and $\phi\to\bot\Rightarrow\neg\phi$ are all forms of RAA, while the rules $\{\neg\phi\to\psi,\neg\phi\to\neg\psi\}\Rightarrow\phi$, $\neg\phi\to\phi\Rightarrow\phi$ (consequentia mirabilis) and $\neg\phi\to\bot\Rightarrow\phi$ are forms of proof by contradiction.
Rule 3.4.16 (Disjunctive syllogism).
$\{\phi\vee\psi,\neg\phi\}\Rightarrow\psi.$
Proof.
1 $\phi\vee\psi$ Premise
2 $\neg\phi$ Premise
3 $(\phi\vee\psi)\wedge\neg\phi$ 1,2: Conjunction introduction
4 $(\phi\wedge\neg\phi)\vee(\psi\wedge\neg\phi)$ 3: Distributivity
5 $\phi\wedge\neg\phi\to\psi$ ECQ
6 $\psi\wedge\neg\phi\to\psi$ Conjunction elimination
7 $\psi$ 4,5,6: Proof by exhaustion
Rule 3.4.17 (Implication and inconsistency).
$\phi\to\psi\Leftrightarrow\phi\wedge\neg\psi\to\bot.$
Proof.
$(\phi\to\psi)$
1     $\leftrightarrow(\phi\wedge\neg\psi\leftrightarrow\bot)$ Definition of implication
2     $\leftrightarrow(\phi\wedge\neg\psi\to\bot)$ Univariate implication expressions
3 $(\phi\to\psi)\leftrightarrow(\phi\wedge\neg\psi\to\bot)$ 1,2: Transitivity
Rule 3.4.18 (LEM exhaustion for implication).
$\{\chi\wedge\phi\to\psi,\neg\chi\wedge\phi\to\psi\}\Rightarrow\phi\to\psi.$
Proof.
1 $\chi\wedge\phi\to\psi$ Premise
2 $\neg\chi\wedge\phi\to\psi$ Premise
3 $(\chi\wedge\phi)\vee(\neg\chi\wedge\phi)\to\psi$ 1,2: Two-to-one
$(\chi\wedge\phi)\vee(\neg\chi\wedge\phi)$
4     $\leftrightarrow(\chi\vee\neg\chi)\wedge\phi$ Distributivity
5     $\leftrightarrow\phi$ Theorem (LEM) like top
6 $(\chi\wedge\phi)\vee(\neg\chi\wedge\phi)\leftrightarrow\phi$ 4–5: Transitivity
7 $\phi\to\psi$ 6,3: Replacement
Rule 3.4.19 (Resolution).
$\{\phi\vee\chi,\psi\vee\neg\chi\}\Rightarrow\phi\vee\psi.$
Proof.
1 $\neg\chi\wedge(\phi\vee\chi)\to\phi$ Disjunctive syllogism
2 $\chi\wedge(\psi\vee\neg\chi)\to\psi$ Disjunctive syllogism
3 $\neg\chi\wedge(\phi\vee\chi)\to\phi\vee\psi$ 1: Strengthening
4 $\chi\wedge(\psi\vee\neg\chi)\to\phi\vee\psi$ 2: Strengthening
5 $\neg\chi\wedge(\phi\vee\chi)\wedge(\psi\vee\neg\chi)\to\phi\vee\psi$ 3: Weakening
6 $\chi\wedge(\phi\vee\chi)\wedge(\psi\vee\neg\chi)\to\phi\vee\psi$ 4: Weakening
7 $(\phi\vee\chi)\wedge(\psi\vee\neg\chi)\to\phi\vee\psi$ 5,6: LEM exhaustion for implication
Rule 3.4.20 (Cut).
$\{\phi_{1}\to\chi,\phi_{2}\wedge\chi\to\psi\}\Rightarrow\phi_{1}\wedge\phi_{2}\to\psi.$
Proof.
1 $\phi_{1}\to\chi$ Premise
2 $\phi_{2}\wedge\chi\to\psi$ Premise
3 $\phi_{1}\wedge\phi_{2}\to\chi$ 1: Weakening
4 $\phi_{1}\wedge\phi_{2}\to\phi_{1}\wedge\phi_{2}\wedge\chi$ 3: Strengthening
5 $\phi_{1}\wedge\phi_{2}\wedge\chi\to\psi$ 2: Weakening
6 $\phi_{1}\wedge\phi_{2}\to\psi$ 4,5: Transitivity
Meta Rule 3.4.1 (Equivalence of implication theorem and disjunction theorem).
$(\vdash\phi\to\psi)\Leftrightarrow(\vdash\neg\phi\vee\psi).$
Proof.
1 $\vdash\phi\to\psi$ Premise
2     $(\phi\to\psi)\to\neg\phi\vee\psi$ Implication-to-Disjunction
3     $\phi\to\psi$ 1 (assumed theorem)
4     $\neg\phi\vee\psi$ 2,3: Modus ponens
5 $\vdash\neg\phi\vee\psi$ 2–4: Unconditional proof
1 $\vdash\neg\phi\vee\psi$ Premise
2     $\phi$ Premise
3     $\neg\phi\vee\psi$ 1 (assumed theorem)
4     $\psi$ 2,3: Disjunctive syllogism
5 $\vdash\phi\to\psi$ 3–5: Conditional proof
4 Semantics and the algebra IRLA
4.1 Semantics of IRL
An interpretation of a formal language assigns a meaning (semantic value) to each expression within that language. However, logic is fundamentally concerned with truth values rather than specific contents of statements of a language. So interpretations that assign the same truth value (even if they assign different contents) to a statement can be viewed as the same in a logic. Hence, an interpretation of a propositional language can be represented by just a truth assignment or valuation function.
Let $\{\mathrm{T,F}\}$ be the set of truth values for a two-valued logic.
Let $L$ be the set of all formulas in the language of IRL.
Let $U=\{\mathrm{T,F}\}^{L}$, the set of all valuation functions for $L$.
The set $L$ is countably infinite (even in a case when the countable set of primitive propositions is finite), thus $|U|=|\{\mathrm{T,F}\}|^{|L|}=2^{|L|}=2^{\aleph_{0}}=\mathfrak{c}$, i.e. $U$ is generally uncountable with the cardinality of $\mathbb{R}$. (In classical logic, as all the operations – including particularly the implication – are truth-functional, the number of all interpretations for a formula with $n$ propositional variables is only $2^{n}$.)
Definition 4.1.1 (T-set).
The T-set of a formula $\phi\in L$ is defined as
$V_{\phi}=\{v\in U\mid v(\phi)=\mathrm{T}\}.$
Semantics of IRL
The semantic system for IRL is defined via T-set as follows.
  1. 1.
    The structure $(2^{U},U,\emptyset,{}^{\complement},\cap,\cup,\subseteq)$ is the power set Boolean lattice of $U$. (For any formula $\phi$ its T-set $V_{\phi}$ is a member of $2^{U}$ by definition.)
  2. 2.
    For any primitive proposition $A$: $V_{A}$ is given (may be unknown, so underivable).
  3. 3.
[…]
    $V_{\top}=U$, $V_{\bot}=\emptyset$.
    $V_{\neg\phi}=V_{\phi}^{\complement}$.
    $V_{\phi\wedge\psi}=V_{\phi}\cap V_{\psi}$, $V_{\phi\vee\psi}=V_{\phi}\cup V_{\psi}$.
  4. 4.
    For the non-truth-functional bi-implication $\leftrightarrow$:
    $V_{\phi\leftrightarrow\phi}=U$, $V_{\phi\leftrightarrow\psi}=V_{\psi\leftrightarrow\phi}$, $V_{\phi\leftrightarrow\psi}\cap V_{\psi\leftrightarrow\chi}\subseteq V_{\phi\leftrightarrow\chi}$;
    $V_{\phi\leftrightarrow\psi}\cap V_{\chi}\subseteq V_{\chi[\phi\mapsto\psi]}$;
    $V_{\phi\leftrightarrow\psi}=U$ if and only if $V_{\phi}=V_{\psi}$.
  5. 5.
    For the non-truth-functional implication $\to$: $V_{\phi\to\psi}=V_{\phi\wedge\psi\leftrightarrow\phi}$.
Remark 4.1.1.
  • Semantics of bi-implication, as it is not truth-functional, is given by a collection of characterizing conditions. Implication is defined by bi-implication.
  • A derived result for implication is $V_{\phi\to\psi}=U\text{ if and only if }V_{\phi}\subseteq V_{\psi}$. In terms of the original meaning of implication, this indicates that $V_{\phi}\subseteq V_{\psi}$ is just the “certain mechanism” of the tautological implication from $\phi$ to $\psi$ (the scope of this mechanism is $U$).
  • By definition of T-set, for any formula $\phi$, $\emptyset\subseteq V_{\phi}\subseteq U$, thus $V_{\bot\to\psi}=U$ and $V_{\phi\to\top}=U$, indicating that ECQ and its dual have “certain mechanisms” of tautological implications, so they do not conflict with the original meaning of implication.
Definition 4.1.2 (Validity and semantic consequence).
With respect to the semantic system, validity/tautology, semantic consequence and equivalence, are defined below:
  • A formula $\phi$ is valid or is a tautology, denoted $\vDash\phi$, if and only if $V_{\phi}=U$.
  • A formula $\psi$ is a semantic consequence of a formula $\phi$, denoted $\phi\vDash\psi$, if and only if $V_{\phi}\subseteq V_{\psi}$.
  • Formulas $\phi$ and $\psi$ are semantically equivalent if $\phi\vDash\psi$ and $\psi\vDash\phi$, or equivalently $V_{\phi}=V_{\psi}$.
The semantic system of IRL defined in this subsection can be viewed as an extension of the power set Boolean lattice of $U$, $(2^{U},U,\emptyset,{}^{\complement},\cap,\cup,\subseteq)$, by adding semantic characteristics of propositional formulas of IRL in the T-set notation. The algebraic system IRLA developed in the next subsection, an extension of the Boolean lattice $(X,1,0,\neg,\wedge,\vee,\leq)$, is simply a “concise version” of the semantic system via T-set in this subsection. Symbols of IRLA language are defined in Table 2.
Table 2: Definition of IRLA symbols
IRL T-set IRLA
Primitive propositions: Constants:
$A,B,...$ $V_{A},V_{B},...$ $a,b,...$
Formulas: Variables:
$\phi,\psi,...$ $V_{\phi},V_{\psi},...$ $x,y,...$
Formula schema: Term:
$F(\phi,\psi,...)$ $V_{F(\phi,\psi,...)}$ $t(x,y,...)$
Operations other than $\leftrightarrow$,$\to$ , →:
,,¬,$\top$,$\bot$,$\neg$,$\wedge$,$\vee$ , ⊥ , ¬ , ∧ , ∨ $U,\emptyset,\cdot^{\complement},\cap,\cup$ $1,0,\neg,\wedge,\vee$
Bi-implication:
$\phi\leftrightarrow\psi$ $V_{\phi\leftrightarrow\psi}$ a $x\leftrightarrow y$ b
Implication:
$\phi\to\psi$ $V_{\phi\to\psi}$ a $x\to y$ b
Tautology:
$\vDash\phi$ $V_{\phi}=U$ $x=1$
Contradiction:
$\vDash\neg\phi$ $V_{\phi}=\emptyset$ $x=0$
Contingency:
$\nvDash\phi$ and $\nvDash\neg\phi$ $V_{\phi}\neq U$ and $V_{\phi}\neq\emptyset$ $x\neq 1$ and $x\neq 0$
Tautological bi-implication:
$\vDash\phi\leftrightarrow\psi$ $V_{\phi\leftrightarrow\psi}=U$ i.e. $V_{\phi}=V_{\psi}$ $x\leftrightarrow y=1$ i.e. $x=y$
Tautological implication:
$\vDash\phi\to\psi$ $V_{\phi\to\psi}=U$ i.e. $V_{\phi}\subseteq V_{\psi}$ $x\to y=1$ i.e. $x\leq y$
  • a
    $V_{\phi\leftrightarrow\psi}$ and $V_{\phi\to\psi}$ can not be generally represented by $V_{\phi}$ and $V_{\psi}$.
  • b
    The IRLA representation of the T-sets $x=V_{\phi}$, $y=V_{\psi}$ and $x\leftrightarrow y=V_{\phi\leftrightarrow\psi}$ implies the definition $V_{\phi}\leftrightarrow V_{\psi}=V_{\phi\leftrightarrow\psi}$, and similarly $V_{\phi}\to V_{\psi}=V_{\phi\to\psi}$.
4.2 Algebra IRLA
The algebra presented in this subsection, called IRLA, is an extension of Boolean lattice that is an extension of Boolean algebra.
Let $(X,1,0,\neg,\wedge,\vee)$ be an algebraic system, where $X$ is the underlying set, $(1,0,\neg,\wedge,\vee)$ are operations of arities $(0,0,1,2,2)$ on $X$. Note that “an operation on a set” implies that the set is closed under the operation. So $1,0\in X$, i.e. $X$ contains at least two constants $1$ and 00. And $X$ may contains other countably many constants $a,b,...$.
Let $x,y,...$ be variables and $s,t,...$ terms for the algebraic language.
Boolean algebra
A Boolean algebra $(X,1,0,\neg,\wedge,\vee)$ is defined by the following axioms and rules of inference.
  • Axioms: commutativity, associativity and distributivity for $\wedge$ and $\vee$; definitions of identity elements $1$ and 00; definition of complement or negation $\neg$. Associativity is redundant (Huntington (1904)).
  • Rules of inference:
    Primitive:
    1. 1.
      $s=s$. (Reflexivity)
    2. 2.
      $\{s=t,u=v\}\Rightarrow(u=v)[s\mapsto t]$. (Replacement property)
    Derived:
    1. 4.
      $s=t\Leftrightarrow t=x$. (Symmetry)
    2. 5.
      $\{s=t,t=u\}\Rightarrow s=u$. (Transitivity)
    3. 6.
      $s=t\Rightarrow u=u[s\mapsto t]$. (Leibniz’s replacement)
Although the above rules are actually default for (a variety of) algebras such as Boolean algebra – so they are usually not listed explicitly, it is best to specify the set of logical rules especially for an algebraic system that represents a logical system (e.g. Boolean algebra and Boolean lattice), even if these rules may be used tacitly.
Boolean lattice
A Boolean lattice $(X,1,0,\neg,\wedge,\vee,\leq)$ is an extension of Boolean algebra $(X,1,0,\neg,\wedge,\vee)$ by adding the defining axiom/rule for the partial order ($\leq$): $x\leq y\Leftrightarrow x\wedge y=x$. The following can be derived:
  1. 1.
    $s\leq s$.
  2. 2.
    $\{s\leq t,t\leq s\}\Rightarrow s=t$.
  3. 3.
    $\{s\leq t,t\leq u\}\Rightarrow s\leq u$, $\{s=t,t\leq u\}\Rightarrow s\leq u$, $\{s\leq t,t=u\}\Rightarrow s\leq u$.
  4. 4.
    $\{s=t,u\leq v\}\Rightarrow(u\leq v)[s\mapsto t]$.
  5. 5.
    $s=t\Rightarrow u[s\mapsto t]\leq u$, $s=t\Rightarrow u\leq u[s\mapsto t]$.
Definition of IRLA
Definition 4.2.1 (Algebra IRLA).
The algebraic system $IRLA=(X,1,0,\neg,\wedge,\vee,\leftrightarrow,\to,\leq)$ is the extension of Boolean lattice $BL=(X,1,0,\neg,\wedge,\vee,\leq)$ by defining the new operations ($\leftrightarrow$ and $\to$) with the axioms below:
  1. 1.
    $(x\leftrightarrow y)\wedge z\leq z[x\mapsto y]$. (Replacement property)
  2. 2.
    $x=y\Rightarrow x\leftrightarrow y=1$. (Tautological bi-implication)
  3. 3.
    $x\to y=x\wedge y\leftrightarrow x$. (Definition of implication)
4.3 Some theorems of IRLA
Theorem 4.3.1 (Reflexivity of bi-implication in IRLA).
$x\leftrightarrow x=1$.
Proof.
1 $x=x$ Reflexivity of $=$
2 $x\leftrightarrow x=1$ 1: Tautological bi-implication (Axiom 2)
Theorem 4.3.2 (Symmetry of bi-implication in IRLA).
$x\leftrightarrow y=y\leftrightarrow x$.
Proof.
1 $x\leftrightarrow x=1$ Reflexivity of $\leftrightarrow$
$x\leftrightarrow y$
2     $=(x\leftrightarrow y)\wedge 1$ Identity element
3     $=(x\leftrightarrow y)\wedge(x\leftrightarrow x)$ 1: Leibniz’s replacement of $=$
4     $\leq(x\leftrightarrow x)[x\mapsto y]$ Replacement property of $\leftrightarrow$ (Axiom 1)
5     $=x[x\mapsto y]\leftrightarrow x$ Selected replacement
6     $=y\leftrightarrow x$ Result of the replacement
7 $x\leftrightarrow y\leq y\leftrightarrow x$ 2–6: Transitivity of $=$ and $\leq$
8 $y\leftrightarrow x\leq x\leftrightarrow y$ 1–7: Similarly
9 $x\leftrightarrow y=y\leftrightarrow x$ 7,8: Antisymmetry of $\leq$
Theorem 4.3.3 (Transitivity of bi-implication in IRLA).
$(x\leftrightarrow y)\wedge(y\leftrightarrow z)\leq x\leftrightarrow z.$
Proof.
$(x\leftrightarrow y)\wedge(y\leftrightarrow z)$
1     $=(y\leftrightarrow z)\wedge(x\leftrightarrow y)$ Commutativity of $\wedge$
2     $\leq(x\leftrightarrow y)[y\mapsto z]$ Replacement property of $\leftrightarrow$ (Axiom 1)
3     $=(x\leftrightarrow z)$ Result of the replacement
4 $(x\leftrightarrow y)\wedge(y\leftrightarrow z)\leq x\leftrightarrow z$ 1–3: Transitivity of $=$ and $\leq$
Theorem 4.3.4 (Leibniz’s replacement in IRLA).
$x\leftrightarrow y\leq z\leftrightarrow z[x\mapsto y].$
Proof.
1 $z\leftrightarrow z=1$ Reflexivity
$x\leftrightarrow y$
2     $=(x\leftrightarrow y)\wedge 1$ Identity element
3     $=(x\leftrightarrow y)\wedge(z\leftrightarrow z)$ 1: Replacement property of $=$
4     $\leq(z\leftrightarrow z)[x\mapsto y]$ Replacement property of $\leftrightarrow$ (Axiom 1)
5     $=z\leftrightarrow z[x\mapsto y]$ Selected replacement
6 $x\leftrightarrow y\leq z\leftrightarrow z[x\mapsto y]$ 2–5: Transitivity of $=$ and $\leq$
Theorem 4.3.5 (Tautological bi-implication and equality).
$x\leftrightarrow y=1\Leftrightarrow x=y.$
Proof.
1 $x\leftrightarrow y=1$ Premise
$x$
2     $=1\wedge x$ Identity element
3     $=(x\leftrightarrow y)\wedge x$ 1: Replacement property of $=$
4     $\leq y$ Replacement property of $\leftrightarrow$ (Axiom 1)
5 $x\leq y$ 2–4: Transitivity of $=$ and $\leq$
6 $y\leftrightarrow x=x\leftrightarrow y$ Symmetry of $\leftrightarrow$
7 $y\leftrightarrow x=1$ 6,1:Transitivity of $=$
8 $y\leq x$ 7,1–5: Similarly
9 $x=y$ 5,8: Antisymmetry of $\leq$
Thus we have $x\leftrightarrow y=1\Rightarrow x=y$, and combining this with $x=y\Rightarrow x\leftrightarrow y=1$ (Axiom 2) we get $x\leftrightarrow y=1\Leftrightarrow x=y$
Corollary 4.3.1 (Tautological implication and inequality).
$x\to y=1\Leftrightarrow x\leq y.$
Proof.
By definition of $\to$ (Axiom 3), $x\to y=1\Leftrightarrow x\wedge y\leftrightarrow x=1$, and $\leq$ is defined by $x\leq y\Leftrightarrow x\wedge y=x$. In terms of the above theorem, $x\wedge y\leftrightarrow x=1\Leftrightarrow x\wedge y=x$, so $x\to y=1\Leftrightarrow x\leq y$. ∎
In terms of Theorem 4.3.5 and Corollary 4.3.1, an IRLA theorem of the form $s\leftrightarrow t=1$ or $s\to t=1$ follows immediately from a corresponding theorem of Boolean lattice (BL) of the form $s=t$ or $s\leq t$, so the latter is also a theorem in IRLA. Some examples of such theorems of IRLA and BL are listed in Table 3.
Table 3: Some theorems of IRLA and BL
Name Form only in IRLA Form in BL and IRLA
Idempotence of $\wedge$ $x\wedge x\leftrightarrow x=1$ $x\wedge x=x$
Idempotence of $\vee$ $x\vee x\leftrightarrow x=1$ $x\vee x=x$
Commutativity of $\wedge$ $x\wedge y\leftrightarrow y\wedge x=1$ $x\wedge y=y\wedge x$
Commutativity of $\vee$ $x\vee y\leftrightarrow y\vee x=1$ $x\vee y=y\vee x$
De Morgan’s law 1 $\neg(x\wedge y)\leftrightarrow\neg x\vee\neg y=1$$\neg(x\wedge y)=\neg x\vee\neg y$
De Morgan’s law 2 $\neg(x\vee y)\leftrightarrow\neg x\wedge\neg y=1$$\neg(x\vee y)=\neg x\wedge\neg y$
Reflexivity of implication $x\to x=1$ $x\leq x$
Conjunction elimination $x\wedge y\to x=1$ $x\wedge y\leq x$
Disjunction introduction $x\to x\vee y=1$ $x\leq x\vee y$
Boundedness of $\to$ $10\to x=1,\;x\to 1=$ 1 $0\leq x,\;x\leq 1$
Theorem 4.3.6 (Bi-implication introduction / Antisymmetry of implication in IRLA).
$(x\to y)\wedge(y\to x)\leq x\leftrightarrow y.$
Proof.
$(x\to y)\wedge(y\to x)$
1     $=(x\leftrightarrow x\wedge y)\wedge(x\wedge y\leftrightarrow y)$ Definition of $\to$
2     $\leq x\leftrightarrow y$ Transitivity of $\leftrightarrow$
3 $(x\to y)\wedge(y\to x)\leq x\leftrightarrow y$ 1–2: Transitivity of $=$ and $\leq$
Theorem 4.3.7 (Bi-implication elimination in IRLA).
$x\leftrightarrow y\leq x\to y,\>x\leftrightarrow y\leq y\to x.$
Proof.
$x\leftrightarrow y$
1     $=(x\leftrightarrow y)\wedge 1$ Identity element
2     $=(x\leftrightarrow y)\wedge(x\to x)$ Reflexivity of $\to$ ($x\leftrightarrow x=1$)
3     $\leq(x\to x)[x\mapsto y]$ Replacement property of $\leftrightarrow$ (Axiom 1)
4     $=x\to x[x\mapsto y]$ Selected replacement
5     $=x\to y$ Result of the replacement
6 $x\leftrightarrow y\leq x\to y$ 1–5: Transitivity of $=$ and $\leq$
The second is obtained by symmetry of $\leftrightarrow$. ∎
Theorem 4.3.8 (Bi-implication by implication in IRLA).
$x\leftrightarrow y=(x\to y)\wedge(y\to x).$
Proof.
1 $x\leftrightarrow y\leq x\to y$ Bi-implication elimination
2 $x\leftrightarrow y\leq y\to x$ Bi-implication elimination
3 $x\leftrightarrow y\leq(x\to y)\wedge(y\to x)$ 1,2: One-to-two of $\leq$
4 $(x\to y)\wedge(y\to x)\leq x\leftrightarrow y$ Bi-implication introduction
5 $x\leftrightarrow y=(x\to y)\wedge(y\to x)$ 3,4: Antisymmetry of $\leq$
Theorem 4.3.9 (Transitivity of implication in IRLA).
$(x\to y)\wedge(y\to z)\leq x\to z.$
Proof.
$(x\to y)\wedge(y\to z)$
1     $=(x\wedge y\leftrightarrow x)\wedge(y\wedge z\leftrightarrow y)$ Definition of $\to$
2     $=(x\wedge y\leftrightarrow x)\wedge(x\wedge y\leftrightarrow x)\wedge(y\wedge z\leftrightarrow y)$ Idempotence
3     $\leq(x\wedge y\leftrightarrow x)\wedge(x\wedge(y\wedge z)\leftrightarrow x)$ Replacement $[y\mapsto y\wedge z]$
4     $=(x\wedge y\leftrightarrow x)\wedge((x\wedge y)\wedge z\leftrightarrow x)$ Associativity
5     $\leq x\wedge z\leftrightarrow x$ Replacement $[x\wedge y\mapsto x]$
6     $=x\to z$ Definition of $\to$
7 $(x\to y)\wedge(y\to z)\leq x\to z$ 1–6: Transitivity of $=$ and $\leq$
Theorem 4.3.10 (One-to-two and Two-to-one in IRLA).
$(z\to x)\wedge(z\to y)$ $=z\to x\wedge y.$
$(x\to z)\wedge(y\to z)$ $=x\vee y\to z.$
Proof.
$(z\to x)\wedge(z\to y)$
1     $=(x\wedge z\leftrightarrow z)\wedge(y\wedge z\to z)$ Definition of $\to$
2     $\leq x\wedge y\wedge z\leftrightarrow z$ Replacement $[z\mapsto y\wedge z]$
3     $=z\to x\wedge y$ Definition of $\to$
4 $(z\to x)\wedge(z\to y)\leq z\to x\wedge y$ 1–3: Transitivity of $=$ and $\leq$
$z\to x\wedge y$
5     $=(z\to x\wedge y)\wedge(x\wedge y\to x)$ Conjunction elimination ($x\wedge y\to x=1$)
6     $\leq z\to x$ Transitivity of $\to$
7 $z\to x\wedge y\leq z\to x$ 5–6: Transitivity of $=$ and $\leq$
8 $z\to x\wedge y\leq z\to y$ 5–7: Similarly
9 $z\to x\wedge y\leq(z\to x)\wedge(z\to y)$ 7,8: One-to-two of $\leq$
10 $(z\to x)\wedge(z\to y)=z\to x\wedge y$ 4,9: Antisymmetry of $\leq$
The second is proved dually. ∎
Theorem 4.3.11 (Modus ponens in IRLA).
$(x\to y)\wedge x\leq y.$
Proof.
$(x\to y)\wedge x$
1     $=(x\leftrightarrow x\wedge y)\wedge x$ Definition of $\to$
2     $\leq x\wedge y$ Replacement property of $\leftrightarrow$ ($[x\mapsto x\wedge y]$)
3     $\leq y$ Conjunction elimination
4 $(x\to y)\wedge x\leq y$ 1–3: Transitivity of $=$ and $\leq$
Corollary 4.3.2 (Upper bound of implication).
$x\to y\leq\neg x\vee y.$
Proof.
In BL, $x\wedge y\leq z\Leftrightarrow x\leq\neg y\vee z$, so modus ponens $(x\to y)\wedge x\leq y$ is equivalent to the upper bound of implication $x\to y\leq\neg x\vee y$ in IRLA, hence the latter follows from the above theorem. ∎
Theorem 4.3.12 (Tautological implication and disjunction).
$x\to y=1\Leftrightarrow\neg x\vee y=1.$
Proof.
It follows from $x\to y=1\Leftrightarrow x\leq y$ of IRLA and $x\leq y\Leftrightarrow\neg x\vee y=1$ of BL. ∎
Theorem 4.3.13 (Contradictory implication).
$1\to 0=0$
Proof.
It follows from boundedness $0\leq 1\to 0$ and $1\to 0\leq\neg 1\vee 0=0\vee 0=0$ by Corollary 4.3.2. ∎
4.4 Soundness and completeness
By definition (cf. Subsection 4.1), the T-set of IRL formula schema $F(\phi,\psi,...)$ is the IRLA term $t(x,y,...)$; and $\vDash\phi$ is the same as $x=1$, $\vDash\phi\leftrightarrow\psi$ is the same as $x=y$ or $x\leftrightarrow y=1$, and $\vDash\phi\to\psi$ is the same as $x\leq y$ or $x\to y=1$.
Soundness of IRL
Meta Theorem 4.4.1 (Axioms are tautologies).
All axioms of IRL are tautologies.
Sketch of proof.
For each axiom schema $F(\phi,\psi,...)$ of IRL, verify $t(x,y,...)=1$ in IRLA. For example, $\phi\leftrightarrow\phi$ is a tautology since $x\leftrightarrow x=1$ in IRLA. ∎
Meta Theorem 4.4.2 (Primitive rules preserve tautologousness).
All the three primitive rules (except for the meta rule of conditional proof) of IRL, i.e. Conjunction introduction $\{\phi,\psi\}\Rightarrow\phi\wedge\psi$, Conjunction elimination $\phi\wedge\psi\Rightarrow\phi$ and Replacement property $\{\phi\leftrightarrow\psi,\chi\}\Rightarrow\chi[\phi\mapsto\psi]$, are tautologousness-preserving, that is, if their premises are tautologies, so are their conclusions.
Proof.
Conjunction introduction $\{\phi,\psi\}\Rightarrow\phi\wedge\psi$:
1 $\vDash\phi$ Known
2 $\vDash\psi$ Known
3 $x=1$ 1: Definition
4 $y=1$ 2: Definition
$x\wedge y$
5     $=1\wedge 1$ 3,4: Replacement
6     $=1$ Identity element
7 $x\wedge y=1$ 5,6: Transitivity
8 $\vDash\phi\wedge\psi$ 7: Definition
Conjunction elimination $\phi\wedge\psi\Rightarrow\phi$:
1 $\vDash\phi\wedge\psi$ Known
2 $x\wedge y=1$ 1: Definition
$1$
3     $=x\wedge y$ 2: Symmetry
4     $\leq x$ Conjunction elimination
5 $1\leq x$ 3,4: Transitivity of $=$ and $\leq$
6 $x\leq 1$ Boundedness
7 $x=1$ 5,6: Antisymmetry of $\leq$
8 $\vDash\phi$ 7: Definition
Replacement property $\{\phi\leftrightarrow\psi,\chi\}\Rightarrow\chi[\phi\mapsto\psi]$:
1 $\vDash\phi\leftrightarrow\psi$ Known
2 $\vDash\chi$ Known
3 $x\leftrightarrow y=1$ 1: Definition
4 $z=1$ 2: Definition
$1$
5     $=1\wedge 1$ Identity element
6     $=(x\leftrightarrow y)\wedge z$ 3,4: Replacement
7     $\leq z[x\mapsto y]$ Replacement property of $\leftrightarrow$
8     $\leq 1$ Boundedness
9 $z[x\mapsto y]=1$ 5–7–8: Transitivity of $=$ and $\leq$, and Antisymmetry of $\leq$
10 $\vDash\chi[\phi\mapsto\psi]$ 9: Definition
Meta Theorem 4.4.3 (CP rule preserves tautologousness).
Suppose a conditional proof $\phi\vdash\psi$ of IRL satisfies that if a member of its proof sequence is a theorem then it is a tautology, then $\phi\to\psi$ is a tautology. That is, the rule of conditional proof (the CP rule) is tautologousness-preserving.
Proof.
We prove by induction that, for any integer $n\geq 1$, for a conditional proof $\phi\vdash\psi$ of length $n$ satisfying that if a member of its proof sequence is a theorem then it is a tautology (let it be called “theorem-is-tautology” condition below), then $\phi\to\psi$ is a tautology.
Only the three primitive rules need considered in the meta proof: Conjunction introduction $\{\phi,\psi\}\Rightarrow\phi\wedge\psi$, Conjunction elimination $\phi\wedge\psi\Rightarrow\phi$, and Replacement property $\{\phi\leftrightarrow\psi,\chi\}\Rightarrow\chi[\phi\mapsto\psi]$.
Let $\alpha$, $\beta$ and $\gamma$ be IRL formulas that exist in the cases discussed below, and their IRLA counterparts be $p$, $q$ and $r$ respectively.
  1. 1.
    For $n=1$: The proof sequence of $\phi\vdash\psi$ of length $1$ is just $(\psi)$. There are following two cases for $\psi$.
    • Case 1: $\psi$ is a theorem. In this case, $\psi$ is a tautology by the “theorem-is-tautology” condition, so $\phi\to\psi$ must be a tautology as shown below:
      1 $\vDash\psi$ Known
      2 $y=1$ 1: Definition
      3 $x\leq 1$ Boundedness
      4 $x\leq y$ 2,3: Replacement
      5 $\vDash\phi\to\psi$ 4: Definition
    • Case 2: $\psi$ is obtained by a one-premise rule. In this case, $\psi$ can only be obtained by the rule $\psi\wedge\alpha\Rightarrow\psi$ where $\psi\wedge\alpha=\phi$, so $\phi\to\psi$ must be a tautology as shown below:
      1 $\psi\wedge\alpha=\phi$ Known
      2 $y\wedge p=x$ 1: Definition
      3 $y\wedge p\leq y$ Conjunction elimination
      4 $x\leq y$ 2,3: Replacement
      5 $\vDash\phi\to\psi$ 4: Definition
    Thus, if $\phi\vdash\psi$ of length $1$ satisfies the “theorem-is-tautology” condition, then $\phi\to\psi$ is a tautology.
  2. 2.
    For any $n\geq 1$: Suppose that, for any $k=1,...,n$, if $\phi\vdash\psi$ of length $k$ satisfies the “theorem-is-tautology” condition, then $\phi\to\psi$ is a tautology; it need to show that if $\phi\vdash\psi$ of length $n+1$ satisfies the “theorem-is-tautology” condition, then $\phi\to\psi$ is a tautology.
    Let $(\psi_{1},...,\psi_{n},\psi)$ be a proof sequence of the length-$(n+1)$ conditional proof $\phi\vdash\psi$. Then, for any $\chi\in\{\phi,\psi_{1},...,\psi_{n}\}$, $\phi\to\chi$ must be a tautology as shown below:
    1 $\chi\in\{\phi,\psi_{1},...,\psi_{n}\}$ Known
    2 $\chi=\phi$ or $\chi\in\{\psi_{1},...,\psi_{n}\}$ 1: Equivalent to it
    3     $\chi=\phi$ Assumption
    4     $z=x$ 3: Definition
    5     $x\leq x$ Reflexivity of $\leq$
    6     $x\leq z$ 4,5: Replacement
    7     $\vDash\phi\to\chi$ 6: Definition
    8     $\chi\in\{\psi_{1},...,\psi_{n}\}$ Assumption
    9     $\vDash\phi\to\chi$ 8: Induction hypothesis (as $\vDash\phi\to\psi_{k}$ for all $k=1,...,n$)
    10 $\vDash\phi\to\chi$ 2,3–7,8–9: Proof by exhaustion
    This result is used in the following four cases for $\psi$.
    • Case 1: $\psi$ is a theorem. In this case, we proceed as for $n=1$.
    • Case 2: $\psi$ is obtained by the rule $\{\alpha,\beta\}\Rightarrow\alpha\wedge\beta$ where $\alpha,\beta\in\{\phi,\psi_{1},...,\psi_{n}\}$ and $\alpha\wedge\beta=\psi$. In this case, $\phi\to\alpha$ and $\phi\to\beta$ are tautologies, so $\phi\to\psi$ must be a tautology as shown below:
      1 $\vDash\phi\to\alpha$ Known
      2 $\vDash\phi\to\beta$ Known
      3 $\alpha\wedge\beta=\psi$ Known
      4 $x\leq p$ 1: Definition
      5 $x\leq q$ 2: Definition
      6 $p\wedge q=y$ 3: Definition
      7 $x\leq p\wedge q$ 4,5: One-to-two
      8 $x\leq y$ 6,7: Replacement
      9 $\vDash\phi\to\psi$ 8: Definition
    • Case 3: $\psi$ is obtained by the rule $\psi\wedge\alpha\Rightarrow\psi$ where $\psi\wedge\alpha\in\{\phi,\psi_{1},...,\psi_{n}\}$. In this case, $\phi\to\psi\wedge\alpha$ is a tautology, so $\phi\to\psi$ must be a tautology as shown below:
      1 $\vDash\phi\to\psi\wedge\alpha$ Known
      2 $x\leq y\wedge p$ 1: Definition
      3 $y\wedge p\leq y$ Conjunction elimination
      4 $x\leq y$ 2,3: Transitivity of $\leq$
      5 $\vDash\phi\to\psi$ 4: Definition
    • Case 4: $\psi$ is obtained by the rule $\{\alpha\leftrightarrow\beta,\gamma\}\Rightarrow\psi$ where $\alpha\leftrightarrow\beta,\gamma\in\{\phi,\psi_{1},...,\psi_{n}\}$ and $\gamma[\alpha\mapsto\beta]=\psi$. In this case, $\phi\to(\alpha\leftrightarrow\beta)$ and $\phi\to\gamma$ are tautologies, so $\phi\to\psi$ must be a tautology as shown below:
      1 $\vDash\phi\to(\alpha\leftrightarrow\beta)$ Known
      2 $\vDash\phi\to\gamma$ Known
      3 $\gamma[\alpha\mapsto\beta]=\psi$ Known
      4 $x\leq p\leftrightarrow q$ 1: Definition
      5 $x\leq r$ 2: Definition
      6 $r[p\mapsto q]=y$ 3: Definition
      7 $x\leq(p\leftrightarrow q)\wedge r$ 4,5: One-to-two
      8 $(p\leftrightarrow q)\wedge r\leq r[p\mapsto q]$ Replacement property of $\leftrightarrow$
      9 $x\leq r[p\mapsto q]$ 7,8: Transitivity of $\leq$
      10 $x\leq y$ 6,9: Replacement
      11 $\vDash\phi\to\psi$ 10: Definition
    Thus, for all $n\geq 1$, suppose that for all $k=1,...,n$, any conditional proof $\phi\vdash\psi$ of length $k$ satisfying the “theorem-is-tautology” condition results in tautological $\phi\to\psi$, then any length-$(n+1)$ conditional proof $\phi\vdash\psi$ satisfying the “theorem-is-tautology” condition leads to tautological $\phi\to\psi$ as well.
Therefore, from 1 and 2 by induction, we have that for any $n\geq 1$, if a conditional proof $\phi\vdash\psi$ of length $n$ satisfies the “theorem-is-tautology” condition, then $\phi\to\psi$ is a tautology. ∎
From Meta Theorem 4.4.1, Meta Theorem 4.4.2 and Meta Theorem 4.4.3, it can be concluded that IRL is sound with respect to its semantics specified by IRLA.
Completeness of IRL
In the following two meta theorems, IRLA axiom $x=y\Rightarrow x\leftrightarrow y=1$ (one of the two defining axioms for $\leftrightarrow$) is considered to be a primitive rule of IRLA, and the universal rule $s=s$ (one of the two primitive universal rules in IRLA) is considered to be an axiom of IRLA.
Meta Theorem 4.4.4.
Any axiom in IRLA has a corresponding theorem in IRL.
Sketch of proof.
  1. 1.
    All axioms of IRLA that are inherited from Boolean lattice have their counterparts in IRL. For example, IRLA axiom $x\wedge y=y\wedge x$ corresponds to IRL axiom $\phi\wedge\phi\leftrightarrow\psi\wedge\phi$.
  2. 2.
    The IRLA axiom $(x\leftrightarrow y)\wedge z\leq z[x\mapsto y]$ (one of the two conditions for defining $\leftrightarrow$ in IRLA) corresponds to IRL theorem $(\phi\leftrightarrow\psi)\wedge\chi\to\chi[\phi\mapsto\psi]$ (from the primitive rule $\{\phi\leftrightarrow\psi,\chi\}\Rightarrow\chi[\phi\mapsto\psi]$ of IRL).
  3. 3.
    The IRLA axiom $x\to y=x\wedge y\leftrightarrow x$ (the defining axiom for $\to$) corresponds to IRL axiom $(\phi\to\psi)\leftrightarrow(\phi\wedge\psi\leftrightarrow\phi)$.
  4. 4.
    The universal one $s=s$ corresponds to IRL axiom $\phi\leftrightarrow\phi$.
Meta Theorem 4.4.5.
For all the two primitive rules of IRLA, namely $x=y\Rightarrow x\leftrightarrow y=1$ and $\{s=t,u=v\}\Rightarrow(u=v)[s\mapsto t]$, if their premises have corresponding theorems in IRL, so are their conclusions.
Proof.
  1. 1.
    $x=y\Rightarrow x\leftrightarrow y=1$: In IRL, $\phi\leftrightarrow\psi$ is a theorem if and only if $(\phi\leftrightarrow\psi)\leftrightarrow\top$ is a theorem by the rule of theorem representation. Thus, if the premise $x=y$ has a corresponding theorem $\phi\leftrightarrow\psi$ in IRL, then the conclusion $x\leftrightarrow y=1$ has the counterpart theorem $(\phi\leftrightarrow\psi)\leftrightarrow\top$ in IRL.
  2. 2.
    $\{s=t,u=v\}\Rightarrow(u=v)[s\mapsto t]$: For the premises $s=t$ and $u=v$, if their counterparts $E\leftrightarrow F$ and $G\leftrightarrow H$ are theorems of IRL, then $(G\leftrightarrow H)[E\mapsto F]$, the counterpart in IRL for the conclusion $(u=v)[s\mapsto t]$, must be a theorem in IRL as shown below:
    1 $\vdash E\leftrightarrow F$ Known
    2 $\vdash G\leftrightarrow H$ Known
    3     $E\leftrightarrow F$ 1 (assumed theorem)
    4     $G\leftrightarrow H$ 2 (assumed theorem)
    5     $(G\leftrightarrow H)[E\mapsto F]$ 3,4: Replacement (IRL rule)
    6 $\vdash(G\leftrightarrow H)[E\mapsto F]$ 3–5: Unconditional proof
From Meta Theorem 4.4.4 and Meta Theorem 4.4.5, it can be concluded that IRL is complete with respect to its semantics specified by IRLA.
4.5 Isomorphism between IRL and IRLA
Definition 4.5.1 (Isomorphism between two systems).
Two proof systems $A$ and $B$ (possibly for different formal languages) are isomorphic (to each other), denoted as $A\cong B$, if the following conditions are met:
  1. 1.
    Expressions of the two systems can be translated to each other. Specifically, there is a bijection between expressions of $A$ and expressions of $B$, where “expressions” include not only formulas of object languages but also expressions of meta languages such as that for rules and meta rules.
  2. 2.
    Axioms and primitive rules of $A$ (after translated to expressions of $B$) can be derived in $B$, and vice versa.
Isomorphism between two systems can be equivalently expressed as that they are interpretable in each other in terms of the concept of interpretability (Tarski, 1953).
Two systems $A$ and $B$ have the following properties related to isomorphism:
  1. 1.
    If $A$ and $B$ are isomorphic, then they are equiconsistent and equidecidable.
  2. 2.
    Suppose $B$ is the semantic system of $A$. Then $A$ and $B$ are isomorphic if and only if $A$ is sound and complete with regard to $B$.
Since IRL is sound and complete with respect to IRLA, there is a natural isomorphism between them as shown in Table 4.
Table 4: Isomorphism between IRL and IRLA
IRL $\cong$ IRLA
Primitive propositions $A,B,...$ $a,b,...$ Constants
Formulas $\phi,\psi,...$ $x,y,...$ Variables
Formula schema $F(\phi,\psi,...)$ $t(x,y,...)$ Term
Operations ,,¬,,,$\top$,$\bot$,$\neg$,$\wedge$,$\vee$,$\leftrightarrow$,$\to$ , ⊥ , ¬ , ∧ , ∨ , ↔ , → $1,0,\neg,\wedge,\vee,\leftrightarrow,\to$
Theorem $\vdash\phi$ $x=1$ Tautology
Theorem of negation $\vdash\neg\phi$ $x=0$ Contradiction
Independent formula $\nvdash\phi$ and $\nvdash\neg\phi$ $x\neq 1$ and $x\neq 0$ Contingency
Bi-implication theorem $\vdash\phi\leftrightarrow\psi$ or $\phi\Leftrightarrow\psi$ a $x\leftrightarrow y=1$ or $x=y$ a Tautological bi-implication
Implication theorem $\vdash\phi\to\psi$ or $\phi\vdash\psi$ or $\phi\Rightarrow\psi$ b $x\to y=1$ or $x\leq y$ b Tautological implication
Meta proof (e.g.)  $(\phi\vdash\psi)\vdash(\vdash\phi\to\psi)$ c $x\leq y\vdash x\to y=1$ c  Conditional proof in IRLA (e.g.)
Meta Rule (e.g.)  $(\phi\vdash\psi)\Rightarrow(\vdash\phi\to\psi)$ c $x\leq y\Rightarrow x\to y=1$ c  Rule in IRLA (e.g.)
  • a
    In IRL, $\vdash\phi\leftrightarrow\psi$ and $\phi\Leftrightarrow\psi$ are equivalent; in IRLA, $x\leftrightarrow y=1$ and $x=y$ are equivalent; so any of the two in IRL can be translated to any of the two in IRLA, and vice versa.
  • b
    In IRL, $\vdash\phi\to\psi$, $\phi\vdash\psi$ and $\phi\Rightarrow\psi$ are equivalent; in IRLA, $x\to y=1$ and $x\leq y$ are equivalent; so any of the three in IRL can be translated to any of the two in IRLA, and vice versa.
  • c
    For simplicity, $\vdash$ are used for both proof and meta proof, and $\Rightarrow$ are used for both rule and meta rule, and they can be used in different systems, if they can be differentiated with context.
As IRL is sound and complete, for a formula $\phi$ of IRL, “$\phi$ is valid or is a tautology” is equivalent to “$\phi$ is a theorem”, “$\phi$ is unsatisfiable or is a contradiction” is equivalent to “$\neg\phi$ is a theorem”; so “$\phi$ is a contingency” is equivalent to “both $\phi$ and $\neg\phi$ are not theorems” i.e. “$\phi$ is independent of IRL”.
4.6 Truth values of implication
The truth values of an implication $\phi\to\psi$ is represented by its T-set $V_{\phi\to\psi}$ i.e. $x\to y$ in IRLA.
Define $x<y$ as $x\leq y$ and $x\neq y$. Thus $10<x<$ 1 means that $x$ is (represents) a contingency.
By Corollary 4.3.2, in any cases we have the bounds of implication as in
$0\leq x\to y\leq\neg x\vee y,$
and both the lower bound 00 and the upper bound $\neg x\vee y$ of $x\to y$ are sharp, i.e. achievable, since we have $1\to 0=0$ and $x\to x=\neg x\vee x$.
Combined with Theorem 4.3.12, there are three specific cases for the truth values of an implication:
$x\to y=1$ if and only if $\neg x\vee y=1$
$x\to y=0$ if $\neg x\vee y=0$ (i.e. $x=1$ and $y=0$)
$10\leq x\to y\leq\neg x\vee y<$ 1 if $10<\neg x\vee y<$ 1
Thus, generally $V_{\phi\to\psi}$ is a subset of $V_{\neg\phi\vee\psi}$ (including of course $\emptyset$ and $V_{\neg\phi\vee\psi}$). Especially in the case where $\neg\phi\vee\psi$ is contingent, unlike classical logic that always adopts the equality $V_{\phi\to\psi}=V_{\neg\phi\vee\psi}$, $V_{\phi\to\psi}$ can be any subset of $V_{\neg\phi\vee\psi}$, i.e. $V_{\phi\to\psi}\in 2^{V_{\neg\phi\vee\psi}}$, so there are $2^{|V_{\neg\phi\vee\psi}|}$ subsets, of which only one is “correct” for $\phi\to\psi$. The correct $V_{\phi\to\psi}$ is determined by the original meaning of implication – a “certain mechanism” (cf. Subsection 2.2) corresponding to the implication $\phi\to\psi$ considered.
Example 4.6.1.
Let $A$ and $B$ are primitive propositions having 4 possible truth-value combinations, i.e. the set of interpretations $U=\{\mathrm{TT,TF,FT,FF}\}$. Then $\emptyset\subset V_{\neg A\vee B}=\{\mathrm{TT,FT,FF}\}\subset U$, so $\neg A\vee B$ is contingent, therefore $V_{A\to B}\in 2^{V_{\neg A\vee B}}=2^{\{\mathrm{TT,FT,FF}\}}$, thus there are $82^{|\{\mathrm{TT,FT,FF}\}|}=2^{3}=$ 8 possibilities and only one of them is correct for $V_{A\to B}$.
The following is an example that shows which should be the “correct one” for the T-set of implication $\phi\to\psi$ as a subset of the T-set of $\neg\phi\vee\psi$.
Example 4.6.2.
Let $O=$“in the open air”, $P=$“it rains”, and $Q=$“the ground gets wet”. Consider the truth values of $P\to Q=$“if it rains, then the ground gets wet”. Suppose: (a) In the open air, whenever it rains the ground must get wet; (b) “it rains” is not the only cause of “the ground gets wet”, no matter it is in the open air or not, e.g., if someone sprinkles water coincidentally on the ground (or the floor), the ground (or the floor) gets wet as well. Then, we obtain the truth values of $P\to Q$ in all 7 possible cases (there is one impossible case that must be excluded: “in the open air, it rains and the ground does not get wet”), as shown in Table 5.
Table 5: Truth table for Example 4.6.2
      $O$       $P$       $Q$       $\neg P\vee Q$ a       $P\to Q$ b
      1       T       T       T       T       T
      2       T       F       T       T       T
      3       T       F       F       T       T
      4       F       T       T       T       F
      5       F       T       F       F       F
      6       F       F       T       T       F
      7       F       F       F       T       F
  • a
    T-set of $\neg P\vee Q$ is the set $\{1,2,3,4,6,7\}$ of rows, indicating that it is contingent.
  • b
    T-set of $P\to Q$, determined by T-set of $O$ rather than that of $P$ and $Q$, is the set $\{1,2,3\}$ of rows, a proper subset of that of $\neg P\vee Q$.
In this example, it is the certain mechanism of “in the open air” that corresponds to the inevitable implication relation of “if it rains, then the ground gets wet”, and the scope of the mechanism is the T-set of the implication. This is a typical example showing that it is not uncommon that $\neg\phi\vee\psi\Rightarrow\phi\to\psi$ is invalid. The example can explain why the classical logic, with the common “implication” unreasonably defined by $\phi\to\psi\Leftrightarrow\neg\phi\vee\psi$, is doomed to fail frequently in common situations in real world.
5 Relationship between IRL and CL
5.1 Definitions and notations
In the following, view a proof system as the set of its axioms and primitive rules, and use “statement” to mean formula or rule expression of a system; treat an object and the singleton containing the object the same thing.
Definition 5.1.1.
Let $A$ be a system. Let $X$ be a set of statements of the system. Denote with $\vdash_{A}X$ that each member $x$ of $X$ is derivable in $A$ (i.e. $x$ is a theorem if it is a formula or $x$ is a derivable rule if it is a rule expression).
Definition 5.1.2 (Supper-system and sub-system).
Let $A$ and $B$ be systems for a same language.
$A$ is a supper-system of $B$, or $B$ is a sub-system of $A$, if $\vdash_{A}B$.
$A$ and $B$ are (deductively) equivalent if $\vdash_{A}B$ and $\vdash_{B}A$, denoted as $A=B$.
Note: For systems $A$ and $B$ for a same language, $A\cong B$ and $A=B$ are equivalent.
Definition 5.1.3 (System extension).
Let $A$ be a system. Let $X$ be a set of statements.
An extension of $A$ with $X$, denoted as $A+X$, is the system obtained from $A$ by adding each member of $X$ as an axiom or a primitive rule.
Definition 5.1.4 (Provable consequence and equivalence).
Let $A$ be a system. Let $X$ and $Y$ be sets of statements.
Then $Y$ is called a provable consequence of $X$ in $A$, if $\vdash_{A+X}Y$.
$X$ and $Y$ are provably equivalent in $A$, if $\vdash_{A+X}Y$ and $\vdash_{A+Y}X$.
Note that $X\vdash_{A}Y$ is different from $\vdash_{A+X}Y$: the former is just the conditional proof of $Y$ from $X$ in $A$, while the latter is not a conditional proof in $A$ – at most we may define the meta conditional proof $(\vdash_{A}X)\vdash_{A}(\vdash_{A}Y)$ as $\vdash_{A+X}Y$.
Definition 5.1.5 (Provable-equivalence class).
Let $A$ be a system and $x$ be a statement.
The provable-equivalence class of $x$ in $A$ is defined as $[x]_{A}=\{y\mid y\textup{ is provably equivalent to }x\textup{ in }A\}$.
Remark 5.1.1.
In terms of the above definitions: an extension of a system is a super-system of that system; the (deductively) equivalent relation “$=$” between two systems can be used in a usual way (i.e. it has the properties of reflexivity, symmetry, transitivity and replacement) ; the extension operator “$+$” is commutative and associative. Let $A$, $B$ and $C$ be sets of statements, any of which can be a system, then:
  • $A=B+C$ implies $\vdash_{A}B$ and $\vdash_{A}C$;
  • $A=A+B$ is equivalent to $\vdash_{A}B$.
5.2 Two super-systems of IRLA
Definition 5.2.1 (BL0 and BL1).
Let $BL=(X,1,0,\neg,\wedge,\vee,\leq)$ be the Boolean lattice described in Subsection 4.2. $BL0=(X,1,0,\neg,\wedge,\vee,\to,\leftrightarrow,\leq)$ and $BL1=(X,1,0,\neg,\wedge,\vee,\to,\leftrightarrow,\leq)$ are extensions of $BL$ defined as
$BL0$ $=BL+\{x\to y=\mu(x,y)\}\text{ where }\mu(x,y)=\begin{cases}1,\;x\leq y\\ 0,\;x\nleq y\end{cases}\text{ and }$
$BL1$ $=BL+\{x\to y=\neg x\vee y\}$
respectively, with $\leftrightarrow$ defined as usual by $x\leftrightarrow y=(x\to y)\wedge(y\to x)$ for both of them.
Remark 5.2.1 (Logics BL0 and BL1 isomorphic to).
  • BL1 is isomorphic to classical (propositional) logic (CL), i.e. $BL1\cong CL$ in symbols.
  • BL0 is isomorphic to a “contra-classical” logic, denoted as CL0, obtained by translating BL0 to a logic in terms of Table 4, so $BL0\cong CL0$. The logic CL0 is contra-classical because the definitions of implication $x\to y=\mu(x,y)$ in BL0 and $x\to y=\neg x\vee y$ in BL1 are contradictory in the general case (cf. Subsection 5.3), hence $BL0+BL1$ ($CL0+CL$) is inconsistent.
Lemma 5.2.1 (Consistency of BL0 and BL1).
BL0 and BL1 are consistent.
Proof.
  • Consistency of BL0. For any given $(x,y)\in X^{2}$, $\mu(x,y)\in X$ uniquely exists (either $1$ or 00), that is, $(x,y)\mapsto\mu(x,y)$ is (indeed) a binary operation on $X$, thus the new operations $\to$ and $\leftrightarrow$ are well-defined. Therefore, BL0 is an extension of BL by definition, hence a conservative extension of BL, thus BL0 is consistent relative to BL (so they are equiconsistent as BL0 is a super-system of BL).
  • Consistency of BL1. BL1 and CL are equiconsistent since $BL1\cong CL$.
Let
$\Delta=\{(x\leftrightarrow y)\wedge z\leq z[x\mapsto y],\;x=y\Rightarrow x\leftrightarrow y=1,\;x\to y=x\wedge y\leftrightarrow x\},$
then
$IRLA=BL+\Delta$
by definition (cf. Subsection 4.2).
Lemma 5.2.2 (BL0 is a super-system of IRLA).
$\vdash_{BL0}IRLA.$
Proof.
Since $\vdash_{BL0}BL$ and $IRLA=BL+\Delta$, we need only to prove $\vdash_{BL0}\Delta$, which is shown below.
  1. 1.
    Proof of $x=y\Rightarrow x\leftrightarrow y=1$ (and $x\neq y\Rightarrow x\leftrightarrow y=0$) in BL0. By definition of $\to$ and $\leftrightarrow$ in BL0:
    If $x=y$, then $x\leq y$ and $y\leq x$, so $x\to y=1$ and $y\to x=1$, thus $x\leftrightarrow y=(x\to y)\wedge(y\to x)=1\wedge 1=1$.
    If $x\neq y$, then $x\nleq y$ or $y\nleq x$, so $x\to y=0$ or $y\to x=0$, thus, let’s say $x\to y=0$, $x\leftrightarrow y=(x\to y)\wedge(y\to x)=0\wedge(x\to y)=0$.
  2. 2.
    Proof of $x\to y=x\wedge y\leftrightarrow x$ in BL0.
    If $x\leq y$ i.e. $x\wedge y=x$ by definition of $\leq$, then $x\to y=1$ and $x\wedge y\leftrightarrow x=1$, so $x\to y=x\wedge y\leftrightarrow x$.
    If $x\nleq y$ i.e. $x\wedge y\neq x$, then $x\to y=0$ and $x\wedge y\leftrightarrow x=0$, so $x\to y=x\wedge y\leftrightarrow x$.
  3. 3.
    Proof of $(x\leftrightarrow y)\wedge z\leq z[x\mapsto y]$ in BL0.
    If $x=y$, then $(x\leftrightarrow y)\wedge z\leq z[x\mapsto y]$ becomes $1\wedge z\leq z[x\mapsto x]$, i.e. $z\leq z$, so it holds.
    If $x\neq y$, then $(x\leftrightarrow y)\wedge z\leq z[x\mapsto y]$ becomes $0\wedge z\leq z[x\mapsto y]$, i.e. $0\leq z[x\mapsto y]$, so it also holds.
Lemma 5.2.3 (BL1 is a super-system of IRLA).
$\vdash_{BL1}IRLA.$
Proof.
Since $\vdash_{BL1}BL$ and $IRLA=BL+\Delta$, we need only to prove $\vdash_{BL1}\Delta$, which is shown below.
  1. 1.
    Proof of $x=y\Rightarrow x\leftrightarrow y=1$ in BL1. If $x=y$, then $x\leftrightarrow y=x\leftrightarrow x=(x\to x)\wedge(x\to x)=x\to x=\neg x\vee x=1$.
  2. 2.
    Proof of $x\to y=x\wedge y\leftrightarrow x$ in BL1. We have $x\wedge y\leftrightarrow x=(x\wedge y\to x)\wedge(x\to x\wedge y)=(\neg(x\wedge y)\vee x))\wedge(\neg x\vee(x\wedge y))=(\neg x\vee\neg y\vee x)\wedge(\neg x\vee(x\wedge y))=1\wedge(\neg x\vee y)=\neg x\vee y=x\to y$.
  3. 3.
    Proof sketch of $(x\leftrightarrow y)\wedge z\leq z[x\mapsto y]$ in BL1.
[…]
    2 Prove $x\leftrightarrow y\leq z\leftrightarrow z[x\mapsto y]$ from 1 by induction;
    3 Prove $(x\leftrightarrow y)\wedge x\leq y$;
    4 Prove $(x\leftrightarrow y)\wedge z\leq z[x\mapsto y]$ from 2 and 3.
Remark 5.2.2 (CL is a super-system of IRL).
Actually, the idea in the proof of the above Lemma 5.2.3 can be used in proving directly that classical propositional logic CL is a supper-system of IRL. This means that the consistency of IRL can be proved before any derivations.
IRLA-BL0 and IRLA-BL1 relations
Meta Theorem 5.2.1 (IRLA-BL0 and IRLA-BL1 relations).
$IRLA+\{x\nleq y\Rightarrow x\to y=0\}=BL0;$
$IRLA+\{\neg x\vee y\leq x\to y\}=BL1.$
Proof.
By definition, $IRLA=BL+\Delta$, $BL0=BL+\{x\to y=\mu(x,y)\}$ where $\{x\to y=\mu(x,y)\}=\{x\leq y\Rightarrow x\to y=1,\>x\nleq y\Rightarrow x\to y=0\}$, and $BL1=BL+\{x\to y=\neg x\vee y\}$ where $\{x\to y=\neg x\vee y\}=\{x\to y\leq\neg x\vee y,\>\neg x\vee y\leq x\to y\}$.
  1. 1.
    IRLA-BL0 relation. $\vdash_{IRLA}\{x\leq y\Rightarrow x\to y=1\}$, i.e. $IRLA=IRLA+\{x\leq y\Rightarrow x\to y=1\}$; $\vdash_{BL0}\Delta$, i.e. $BL0+\Delta=BL0$. Thus
    $IRLA$ +$+\{x\nleq y\Rightarrow x\to y=0\}$
    $=IRLA+\{x\leq y\Rightarrow x\to y=1\}+\{x\nleq y\Rightarrow x\to y=0\}$
    $=IRLA+\{x\leq y\Rightarrow x\to y=1,\>x\nleq y\Rightarrow x\to y=0\}$
    $=IRLA+\{x\to y=\mu(x,y)\}$
    $=BL+\Delta+\{x\to y=\mu(x,y)\}$
    $=BL0+\Delta$
    $=BL0.$
  2. 2.
    IRLA-BL1 relation. $\vdash_{IRLA}\{x\to y\leq\neg x\vee y\}$, i.e. $IRLA=IRLA+\{x\to y\leq\neg x\vee y\}$; $\vdash_{BL1}\Delta$, i.e. $BL1+\Delta=BL1$. Thus
    $IRLA$ +$+\{\neg x\vee y\leq x\to y\}$
    $=IRLA+\{x\to y\leq\neg x\vee y\}+\{\neg x\vee y\leq x\to y\}$
    $=IRLA+\{x\to y\leq\neg x\vee y,\>\neg x\vee y\leq x\to y\}$
    $=IRLA+\{x\to y=\neg x\vee y\}$
    $=BL+\Delta+\{x\to y=\neg x\vee y\}$
    $=BL1+\Delta$
    $=BL1.$
In contrast, Heyting lattice ($HL$) contains $\neg x\vee y\leq x\to y$ but not $x\to y\leq\neg x\vee y$, as a matter of fact $HL+\{x\to y\leq\neg x\vee y\}=BL1$.
5.3 IRL-CL relationship
Disjunction-to-Implication (D2I) is the formula schema $\neg\phi\vee\psi\to(\phi\to\psi)$ in IRL, whose translation is the term $\neg x\vee y\to(x\to y)$ in IRLA.
From the facts $IRL\cong IRLA$, $CL\cong BL1$ (cf. Subsection 4.5) and $IRLA+\{\neg x\vee y\leq x\to y\}=BL1$ (cf. Subsection 5.2), we have $IRL+\{\neg\phi\vee\psi\to(\phi\to\psi)\}=CL$.
It is shown that $D2I$ is generally independent of IRL in this section. Since $D2I$ is a theorem of the consistent CL, $\neg D2I$ is not derivable in CL, so $\neg D2I$ is not derivable in IRL either, as $\vdash_{CL}IRL$. Thus, it needs only to show that $D2I$ is underivable in IRL.
Disjunction-to-Implication is underivable in IRL
Lemma 5.3.1 (D2I is not provable in BL0).
In $BL0=(X,1,0,\wedge,\vee,\to,\leftrightarrow,\leq)$, if there exists $c\in X$ such that $c\neq 1$ and $c\neq 0$, then $\neg x\vee y\leq x\to y$ is not a theorem of BL0.
Proof.
From $c\neq 1$, $1\nleq c$, so $01\to c=$ 0 by definition; similarly, from $c\neq 0$, $c\to 0=0$. Let $t(x,y)=\neg x\vee y\to(x\to y)$. Then $t(1,c)=\neg 1\vee c\to(1\to c)=0\vee c\to 0=c\to 0=0\neq 1$, so $t(x,y)=1$ is not a theorem of BL0 since its instance $t(1,c)\neq 1$. ∎
Lemma 5.3.2 (Condition for D2I being a theorem of IRLA).
In $IRLA=(X,1,0,\wedge,\vee,\leftrightarrow,\to,\leq)$, $\neg x\vee y\to(x\to y)=1$ is a theorem if and only if there exists no $c\in X$ such that $c\neq 1$ and $c\neq 0$.
Proof.
Let $t(x,y)=\neg x\vee y\to(x\to y)$.
Necessity. If there exists $c\in X$ such that $c\neq 1$ and $c\neq 0$, then $t(x,y)=1$ is not a theorem of BL0 in terms of Lemma 5.3.1, neither is it a theorem of IRLA, since $\vdash_{BL0}IRLA$.
Sufficiency. If there exists no $c\in X$ such that $c\neq 1$ and $c\neq 0$, then for all $x,y\in X$, either $x=1$ or $x=0$ (i.e. $x,y\in\{1,0\}$). In IRLA, $t(1,1)=\neg 1\vee 1\to(1\to 1)=1\to 1=1$, $t(1,0)=\neg 1\vee 0\to(1\to 0)=0\to 0=1$, $t(0,1)=\neg 0\vee 1\to(0\to 1)=1\to 1=1$, and $t(0,0)=\neg 0\vee 0\to(0\to 0)=1\to 1=1$. Thus, for all $x,y\in X$, $\neg x\vee y\to(x\to y)=1$, namely it is a theorem of IRLA in this case. ∎
The general case and the no-contingency case
Propositions are in the three types: tautology, contradiction, and contingency.
There are two cases according to what types of propositional formulas the system contains, as described below.
The general case
The system contains all the three types of propositional formulas.
The no-contingency case
The system contains no contingencies.
Both IRL and CL always contain tautologies (e.g. $\phi\vee\neg\phi$) and contradictions (e.g. $\phi\wedge\neg\phi$). So the only difference between the general case and the no-contingency case is whether or not the system contains a contingency.
[…]
Thus, for IRL and CL, it is the general case if and only if there exists a primitive contingency in the system, and it is the no-contingency case if and only if there exists no primitive contingency i.e. any formula is either tautological or contradictory.
For example, suppose the set of primitive propositional symbols (the signature of propositional language) is simply the singleton $\{A\}$, and let $S_{A}$ be an instance of IRL or CL with this signature. Then: (a) $S_{A}$ contains all the three types of propositions if neither $A$ nor $\neg A$ is an axiom; (b) $S_{A}$ contains no contingencies (each formula is either a tautology or a contradiction) if either $A$ or $\neg A$ is an axiom.
When not specified, we mean it is in the general case.
Summary
Based on the previous results, the logic IRL has the relationship to classical logic (CL) as summarized below.
In any cases (no matter the general case or the no-contingency case):
$IRL+\{\neg\phi\vee\psi\to(\phi\to\psi)\}=CL,$
thus IRL is a sub-system of CL, so IRL is consistent (relative to CL), and the negation of Disjunction-to-Implication ¬($\neg(\neg\phi\vee\psi\to(\phi\to\psi))$ is not a theorem of IRL.
  • In the general case (the system contains all types of propositions including contingencies, which is equivalent to that there is a primitive contingency):
[…]
    $\neg\phi\vee\psi\to(\phi\to\psi)\textup{ is independent of }IRL.$
    So IRL is a proper sub-system of CL in the general case.
  • In the no-contingency case (the system contains no contingencies, which is equivalent to that there is no primitive contingency):
    $IRL=CL\textup{, and }\neg\phi\vee\psi\to(\phi\to\psi)\textup{ is a theorem of the system.}$.
Remark 5.3.1.
In the no-contingency case, any formula is either tautological or contradictory. However, it does not mean that the system is syntactically complete in this case, because we only know that every primitive proposition is either a tautology or a contradiction, but we do not know which one it is, and this is underivable.
5.4 Provable-equivalence class of Disjunction-to-Implication
Meta Theorem 5.4.1 (Some members of the provable-equivalence class of Disjunction-to-Implication).
The following formulas are provably equivalent in IRL.
  1. 1.
    $\phi\to(\top\to\phi)$, $\neg\phi\to(\phi\to\bot)$. (Lower bound of univariate implication)
  2. 2.
    $\phi\to(\psi\to\phi)$, $\neg\phi\to(\phi\to\psi)$. (Generalized ECQ)
  3. 3.
    $\neg\phi\vee\psi\to(\phi\to\psi)$. (Disjunction-to-Implication):
  4. 4.
    $(\phi\wedge\psi\to\chi)\to(\phi\to(\psi\to\chi))$. (Exportation)
  5. 5.
    $(\phi\to(\psi\to\chi))\leftrightarrow(\psi\to(\phi\to\chi))$. (Commutativity of antecedents)
  6. 6.
    $\phi\to((\phi\to\psi)\to\psi)$. (Assertion)
  7. 7.
    $(\phi\to\psi)\vee(\psi\to\phi)$. (Totality)
  8. 8.
    $(\phi\to\psi)\vee(\phi\to\neg\psi)$. (Conditional excluded middle)
  9. 9.
    $(\top\to\phi)\vee(\phi\to\bot)$. (Top-or-bottom)
Note that each of 1 and 2 is a pair of dual formulas that are provably equivalent in IRL.
Proof.
It is done in the structure 1$\to$2$\to$3$\to$4$\to$5$\to$6$\to$1; 2$\to$7$\to$9$\to$1; 2$\to$8$\to$9 as follows.
First group 1$\to$2$\to$3$\to$4$\to$5$\to$6$\to$1:
(1) Lower bound of univariate implication $\to$ (2) Generalized ECQ:
1 $\phi$ Premise 2 $\top\to\phi$ 1: Lower bound of univariate implication 3 $\psi\to\phi$ 2: Safe generalized ECQ
(2) Generalized ECQ $\to$ (3) Disjunction-to-Implication:
1 $\neg\phi\to(\phi\to\psi)$ Generalized ECQ 2 $\psi\to(\phi\to\psi)$ Generalized ECQ 3 $\neg\phi\vee\psi\to(\phi\to\psi)$ 1,2: Two-to-one
(3) Disjunction-to-Implication $\to$ (4) Exportation (with D2I, material implication $\phi\to\psi\Leftrightarrow\neg\phi\vee\psi$ holds):
1 $\phi\wedge\psi\to\chi$ Premise 2 $\neg(\phi\wedge\psi)\vee\chi$ 1: Material implication 3 ($(\neg\phi\vee\neg\psi)\vee\chi$ 2: De Morgan’s law 4 $\neg\phi\vee(\neg\psi\vee\chi)$ 3: Associativity 5 $\neg\phi\vee(\psi\to\chi)$ 4: Material implication 6 $\phi\to(\psi\to\chi)$ 5: Material implication
(4) Exportation $\to$ (5) Commutativity of antecedents:
1 $\phi\to(\psi\to\chi)$ Premise 2 $\phi\wedge\psi\to\chi$ 1: Importation 3 $\psi\wedge\phi\to\chi$ 2: Commutativity 4 $\psi\to(\phi\to\chi)$ 3: Exportation
(5) Commutativity of antecedents $\to$ (6) Assertion:
1 $(\phi\to\psi)\to(\phi\to\psi)$ Reflexivity 2 $\phi\to((\phi\to\psi)\to\psi)$ 1: Commutativity of antecedents
(6) Assertion $\to$ (1) Lower bound of univariate implication:
1 $(\phi\to\phi)\leftrightarrow\top$ Reflexivity 2 $\phi\to((\phi\to\phi)\to\phi)$ Assertion 3 $\phi\to(\top\to\phi)$ 1,2: Replacement
Second group 2$\to$7$\to$9$\to$1:
(2) Generalized ECQ $\to$ (7) Totality:
1 $\psi\to(\phi\to\psi)$ Generalized ECQ 2 $\neg\psi\to(\psi\to\phi)$ Generalized ECQ 3 $\psi\vee\neg\psi\to(\phi\to\psi)\vee(\psi\to\phi)$ 1,2: Monotonicity of $\vee$ w.r.t $\to$ 4 $\psi\vee\neg\psi$ LEM 5 $(\phi\to\psi)\vee(\psi\to\phi)$ 3,4: Modus ponens
(7) Totality $\to$ (9) Top-or-bottom:
1 ($(\neg\phi\to\phi)\leftrightarrow(\top\to\phi)$ Theorem 3.4.19 2 $(\phi\to\neg\phi)\leftrightarrow(\phi\to\bot)$ Theorem 3.4.19 3 ($(\neg\phi\to\phi)\vee(\phi\to\neg\phi)$ Totality 4 $(\top\to\phi)\vee(\phi\to\bot)$ (1,2),3: Replacement
(9) Top-or-bottom $\to$ (1) Lower bound of univariate implication:
1 $\phi$ Premise 2 $(\phi\to\bot)\to\neg\phi$ Upper bound of univariate implication 3 ¬(ϕ)italic-$\neg(\phi\to\bot)$ 1,2: Modus tollens 4 $(\top\to\phi)\vee(\phi\to\bot)$ Top-or-bottom 5 $\top\to\phi$ 3,4: Disjunctive syllogism
Third group 2$\to$8$\to$9:
(2) Generalized ECQ $\to$ (8) Conditional excluded middle:
1 $\psi\to(\phi\to\psi)$ Generalized ECQ 2 $\neg\psi\to(\phi\to\neg\psi)$ Generalized ECQ 3 $\psi\vee\neg\psi\to(\phi\to\psi)\vee(\phi\to\neg\psi)$ 1,2: Monotonicity of $\vee$ w.r.t. $\to$ 4 $\psi\vee\neg\psi$ LEM 5 $(\phi\to\psi)\vee(\phi\to\neg\psi)$ 3,4: Modus ponens
(8) Conditional excluded middle $\to$ (9) Top-or-bottom:
1 ($(\top\to\neg\phi)\leftrightarrow(\phi\to\bot)$ Contraposition, Double negation 2 $(\top\to\phi)\vee(\top\to\neg\phi)$ Conditional excluded middle 3 $(\top\to\phi)\vee(\phi\to\bot)$ 1,2: Replacement
[…]
Meta Theorem 5.4.2 (Conjunction-to-equivalence).
The formula $\phi\wedge\psi\to(\phi\leftrightarrow\psi)$ (Conjunction-to-equivalence) is a member of [$[\neg\phi\vee\psi\to(\phi\to\psi)]_{IRL}$.
Proof.
It is done in the chain Generalized ECQ (known) $\to$ Conjunction-to-equivalence (new) $\to$ Lower bound of univariate implication (known):
Generalized ECQ $\to$ Conjunction-to-equivalence:
1 $\phi\wedge\psi$ Premise 2 $\phi$ 1: Conjunction elimination 3 $\psi\to\phi$ 2: Generalized ECQ 4 $\phi\to\psi$ 1–3: Similarly 5 $\phi\leftrightarrow\psi$ 3,4: Bi-implication introduction
Conjunction-to-equivalence $\to$ Lower bound of univariate implication:
1 $\phi$ Premise 2 $\top\wedge\phi$ 1: Identity element 3 $\top\leftrightarrow\phi$ 2: Conjunction-to-equivalence 4 $\top\to\phi$ 3: Bi-implication elimination
All members of [$[\neg\phi\vee\psi\to(\phi\to\psi)]_{IRL}$ are not provable in IRL, unless the extra assumption of no-contingency is made.
6 The decision problem
Since $IRL\cong IRLA$ (cf. Subsection 4.5), a formula schema $F(\phi_{1},...,\phi_{n})$ in IRL corresponds to the term $t(x_{1},...,x_{n})$ in IRLA (cf. Table 4).
In a sentence $t(x_{1},...,x_{n})=1$ of an algebraic system with an underlying set $X$, all variables are implicitly universally quantified over $X$. So the decision problem considered in $IRLA=(X,1,0,\neg,\wedge,\vee,\leftrightarrow,\to,\leq)$ is, given any term $t(x_{1},...,x_{n})$, decide whether or not $t(x_{1},...,x_{n})=1$ for all $(x_{1},...,x_{n})\in X^{n}$ – we may simply say whether or not “$t(x_{1},...,x_{n})$ is valid” for convenience.
A complete decision procedure is not found for IRL/IRLA in this work. A conjecture: IRL/IRLA is undecidable. If it is indeed undecidable, this is a reflection of the nature of implication relation, which comes from the original meaning of implication.
If $t(x_{1},...,x_{n})$ is valid in IRLA, it is valid in both its super-systems BL0 and BL1 (cf. Subsection 5.2). So if $t(x_{1},...,x_{n})$ is invalid in either BL0 or BL1, it is invalid in IRLA. Thus, BL0 and BL1 can be used as two basic validity “filters” for IRLA. For example, ¬($\neg(\neg x\vee y\to(x\to y))$ is invalid in BL1, so it is invalid in IRLA.
Since BL1 represents classical propositional logic (cf. Subsection 5.2), many ready decision procedures can be utilized, such as the basic methods of truth table, truth tree, and SAT algorithms.
The following is a simple validity-checking method for BL0.
Validity checking in BL0 on a minimal set
Consider the validity-checking problem in $BL0=(X,1,0,\neg,\wedge,\vee,\to,\leq)$: given any expression $t(x_{1},...,x_{n})$, whether or not $t(x_{1},...,x_{n})$ is valid, i.e. if $t(x_{1},...,x_{n})=1$ for all $(x_{1},...,x_{n})\in X^{n}$?
Suppose there exists $c\in X$ such that $10<c<$ 1 (i.e. there exists a contingency in the system). Consider validity checking on the set $\{0,c,1\}$, a minimal subset of $X$ that can reflect the general case (there are all the three types of propositions: tautology, contradiction and contingency).
[…]
Table 6: Closure of the set $\{0,c,1\}$ in BL0 *
      $x$       $y$       $\neg x$       $x\wedge y$       $x\vee y$       $x\to y$
      00       00       $1$       00       00       $1$
      00       $c$       $1$       00       $c$       $1$
      00       $1$       $1$       00       $1$       $1$
      $c$       00       $\neg c$       00       $c$       00
      $c$       $c$       $\neg c$       $c$       $c$       $1$
      $c$       $1$       $\neg c$       $c$       $1$       $1$
      $1$       00       00       00       $1$       00
      $1$       $c$       00       $c$       $1$       00
      $1$       $1$       00       $1$       $1$       $1$
[…]
Thus, in BL0, any expression $t(x_{1},...,x_{n})\in\{0,c,\neg c,1\}$ for all $(x_{1},...,x_{n})\in\{0,c,1\}^{n}$. Note that $\neg c$ is also a contingency, i.e. $10<\neg c<$ 1, since, only in a Boolean algebra, $10<x<$ 1 if and only if $10<\neg x<$ 1 for all $x$ in its underlying set.
If $t(x_{1},...,x_{n})\neq 1$ for some $(x_{1},...,x_{n})\in\{0,c,1\}^{n}$, then $t(x_{1},...,x_{n})\neq 1$ for some $(x_{1},...,x_{n})\in X^{n}$, so the answer to the validity question of $t(x_{1},...,x_{n})$ is “No”.
Otherwise $t(x_{1},...,x_{n})=1$ for all $(x_{1},...,x_{n})\in\{0,c,1\}^{n}$, the answer to the validity question of $t(x_{1},...,x_{n})$ is “Unknown”, i.e., it is possible for $t$ is valid and it is also possible for $t$ is invalid, since, although $t=1$ for all $(x_{1},...,x_{n})\in\{0,c,1\}^{n}$, it is possible that $t\neq 1$ for some $(x_{1},...,x_{n})\in X^{n}\setminus\{0,c,1\}^{n}$.
In terms of these facts, we have a simple algorithm for the (in)validity-checking problem as described below.
Validity checking by exhaustion over $\{0,c,1\}\subset X$
In BL0:
  1. 1.
    Input an expression $t(x_{1},...,x_{n})$;
  2. 2.
    For each $(x_{1},...,x_{n})\in\{0,c,1\}^{n}$, compute $t(x_{1},...,x_{n})$;
  3. 3.
    At any step:
    • If $t(u_{1},...,u_{n})\neq 1$ (i.e. $t=0,c,\neg c$) for some $(u_{1},...,u_{n})\in\{0,c,1\}^{n}$, $t$ is Invalid with a countermodel $(x_{1},...,x_{n})=(u_{1},...,u_{n})$;
    • If $t(x_{1},...,x_{n})\neq 0$ (i.e. $\neg t\neq 1$) for some $(u_{1},...,u_{n})\in\{0,c,1\}^{n}$, $\neg t$ is invalid with a countermodel $(x_{1},...,x_{n})=(u_{1},...,u_{n})$;
    • If both $t$ and $\neg t$ are invalid, $t$ is independent of the system;
  4. 4.
    If $t(x_{1},...,x_{n})=1$ for all $(x_{1},...,x_{n})\in\{0,c,1\}^{n}$, the validity of $t$ is Unkown (but $\neg t$ is invalid);
  5. 5.
    If $t(x_{1},...,x_{n})=0$ for all $(x_{1},...,x_{n})\in\{0,c,1\}^{n}$, $t$ is Invalid (but the validity of $\neg t$ is unknown).
For IRLA as a sub-system of BL0:
  • $t$ is invalid in IRLA if it is invalid in BL0.
  • $t$ is independent of IRLA if it is independent of BL0.
  • The validity of $t$ is unknown in IRLA when its validity is unknown in BL0, or $t$ is valid in BL0.
This simple exhaustive search algorithm, with the worst-case complexity $3^{n}$, is of course incomplete, since it is not for all inputs that the algorithm can give “Yes” or “No” answer (for some inputs the answer is “Unknown”). But it is useful as a basic validity “filter” for IRLA/IRL.
Example 6.0.1.
Is $t(x)=(1\to x)\to x$ valid? For $x\in\{0,c,1\}$:
  • 1.
    $x=0$: $t(0)=(1\to 0)\to 0=0\to 0=1$.
  • 2.
    $x=c$: $t(c)=(1\to c)\to c=0\to c=1$.
  • 3.
    $x=1$: $t(1)=(1\to 1)\to 1=1\to 1=1$.
The result is $t(x)=(1\to x)\to x=1$ for all $x$ in $\{0,c,1\}$, so the answer to the validity question of $(1\to x)\to x$ is “Unknown” (i.e. $(1\to x)\to x$ is possibly valid and possibly invalid as well).
Example 6.0.2.
Is $t(x)=x\to(1\to x)$ valid? For $x\in\{0,c,1\}$:
  • 1.
    $x=0$: $t(0)=0\to(1\to 0)=0\to 0=1$.
  • 2.
    $x=c$: $t(c)=c\to(1\to c)=c\to 0=0$. (a countermodel)
So $t(x)=x\to(1\to x)$ is invalid with a countermodel $x=c$. And since $\neg t(0)=\neg 1=0$, $\neg t(x)$ is also invalid, thus $t(x)$ is independent of IRLA.
Example 6.0.3.
Is $t(x,y)=(x\to y)\to\neg x\vee y$ valid? For $(x,y)\in\{0,c,1\}^{2}$:
  • 1.
    $(x,y)=(0,0)$: $t(0,0)=(0\to 0)\to\neg 0\vee 0=1\to 1=1$.
  • 2.
    $(x,y)=(0,c)$: $t(0,c)=(0\to c)\to\neg 0\vee c=1\to 1=1$.
  • 3.
    $(x,y)=(0,1)$: $t(0,1)=(0\to 1)\to\neg 0\vee 1=1\to 1=1$.
  • 4.
    $(x,y)=(c,0)$: $t(c,0)=(c\to 0)\to\neg c\vee 0=0\to\neg c=1$.
  • 5.
    $(x,y)=(c,c)$: $t(c,c)=(c\to c)\to\neg c\vee c=1\to 1=1$.
  • 6.
    $(x,y)=(c,1)$: $t(c,1)=(c\to 1)\to\neg c\vee 1=1\to 1=1$.
  • 7.
    $(x,y)=(1,0)$: $t(1,0)=(1\to 0)\to\neg 1\vee 0=0\to 0=1$.
  • 8.
    $(x,y)=(1,c)$: $t(1,c)=(1\to c)\to\neg 1\vee c=0\to c=1$.
  • 9.
    $(x,y)=(1,1)$: $t(1,1)=(1\to 1)\to\neg 1\vee 1=1\to 1=1$.
The result is $t(x,y)=(x\to y)\to\neg x\vee y=1$ for all $(x,y)$ in $\{0,c,1\}^{2}$, so the answer to the validity question of $(x\to y)\to\neg x\vee y$ is “Unknown” (i.e. $(x\to y)\to\neg x\vee y$ is possibly valid and possibly invalid as well).
Example 6.0.4.
Is $t(x,y)=\neg x\vee y\to(x\to y)$ valid? For $(x,y)\in\{0,c,1\}^{2}$:
  • 1.
    $(x,y)=(0,0)$: $t(0,0)=\neg 0\vee 0\to(0\to 0)=1\to 1=1$.
  • 2.
    $(x,y)=(0,c)$: $t(0,c)=\neg 0\vee c\to(0\to c)=1\to 1=1$.
  • 3.
    $(x,y)=(0,1)$: $t(0,1)=\neg 0\vee 1\to(0\to 1)=1\to 1=1$.
  • 4.
    $(x,y)=(c,0)$: $t(c,0)=\neg c\vee 0\to(c\to 0)=\neg c\to 0=1\to c=0$. (a countermodel)
So $t(x,y)=\neg x\vee y\to(x\to y)$ is invalid with a countermodel $(x,y)=(c,0)$. And as we also have e.g. $t(0,0)=1$, $t(x,y)$ is independent of IRLA.
Example 6.0.5.
Are $t(x,y)=((x\to y)\to x)\to x$ and its negation $\neg t$ valid? For $(x,y)\in\{0,c,1\}^{2}$:
  • 1.
    $(x,y)=(0,0)$: $t(0,0)=((0\to 0)\to 0)\to 0=(1\to 0)\to 0=0\to 0=1$. (a countermodel for $\neg t$)
  • 2.
    $(x,y)=(0,c)$: $t(0,c)=((0\to c)\to 0)\to 0=(1\to 0)\to 0=0\to 0=1$.
  • 3.
    $(x,y)=(0,1)$: $t(0,1)=((0\to 1)\to 0)\to 0=(1\to 0)\to 0=0\to 0=1$.
  • 4.
    $(x,y)=(c,0)$: $t(c,0)=((c\to 0)\to c)\to c=(0\to c)\to c=1\to c=0$. (a countermodel for $t$)
So, $t(x,y)=((x\to y)\to x)\to x$ is invalid with a countermodel $(x,y)=(c,0)$, and $\neg t(x,y)$ is also invalid with a countermodel $(x,y)=(0,0)$. Thus $((x\to y)\to x)\to x$ is independent of IRLA.
Example 6.0.5 is actually about the validity of Peirce’s law, and the result indicates that Peirce’s law, like Disjunction-to-Implication, is independent of IRL. It can be shown that, in IRL, Peirce’s law is a provable consequence of Disjunction-to-Implication, but (very likely) not vice versa.
7 Natural deduction
The natural deduction for IRL is given by the set of rules below:
1. $(\phi\vdash\psi)\Rightarrow(\vdash\phi\to\psi)$toI CP (Restricted, cf. Remark 7.0.1 below)
2. $\{\phi\to\psi,\phi\}\Rightarrow\psi$toE MP
3. $(\phi\to\psi)\Rightarrow(\psi\to\chi)\to(\phi\to\chi)$ Suffixing
4. $\{\phi,\psi\}\Rightarrow\phi\wedge\psi$wedgeI Conjunction
5. $\phi\wedge\psi\Rightarrow\phi$, $\phi\wedge\psi\Rightarrow\psi$wedgeE Simplification
6. $\{\chi\to\phi,\chi\to\psi\}\Rightarrow\chi\to\phi\wedge\psi$ “One-to-two”
7. $\phi\Rightarrow\phi\vee\psi$, $\psi\Rightarrow\phi\vee\psi$veeI Addition
[…]
9. $\phi\wedge(\psi\vee\chi)\Rightarrow(\phi\wedge\psi)\vee(\phi\wedge\chi)$ Distribution
[…]
14. $\{\phi\to\psi,\psi\to\phi\}\Rightarrow\phi\leftrightarrow\psi$leftrightarrowI Bi-implication introduction
15. $\phi\leftrightarrow\psi\Rightarrow\phi\to\psi$, $\phi\leftrightarrow\psi\Rightarrow\psi\to\phi$leftrightarrowE Bi-implication elimination
Remark 7.0.1 (Restriction on CP).
By the definition of proof in Subsection 3.1, any member of a proof sequence must be either a theorem (or a theorem indicated in an outer layer), or a result from a subset of the premises and previous members in its own layer by a rule (or by a rule indicated in an outer layer), not allowing any other cases. Attention should be paid to this restriction for nested proofs, especially for nested conditional proofs.
Equivalence to Hilbert system
The natural deduction for IRL is equivalent to its Hilbert-style system defined in Section 3.
The primitive rules listed above of the natural deduction are already proved in its Hilbert system (cf. Section 3).
On the other hand, all axioms and primitive rules of the Hilbert-style IRL can be derived in this natural deduction. Some derivations are shown as examples in the following.
Rule 7.0.1 (Commutativity of conjunction).
$\phi\wedge\psi\Rightarrow\psi\wedge\phi.$
Proof.
1     $\phi\wedge\psi$ Premise
[…]
5 $\phi\wedge\psi\vdash\psi\wedge\phi$ 1–4: Definition of conditional proof
6 $\phi\wedge\psi\Rightarrow\psi\wedge\phi$ 5: Definition of derivable rule
The last two lines will be omitted in the subsequent proofs.
Rule 7.0.2 (Associativity of conjunction).
$(\phi\wedge\psi)\wedge\chi\Rightarrow\phi\wedge(\psi\wedge\chi).$
Proof.
1 $(\phi\wedge\psi)\wedge\chi$ Premise
[…]
With the rules of commutativity and associativity of conjunction, conjunction introduction and elimination can be generalized to $\{\phi_{1},...,\phi_{n}\}\Rightarrow\phi_{1}\wedge\cdots\wedge\phi_{n}$ and $\phi_{1}\wedge\cdots\wedge\phi_{n}\Rightarrow\phi_{i}$ for any $n\geq 2$, respectively.
Meta Rule 7.0.1 (Equivalence of the three forms).
$(\vdash\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi)\Leftrightarrow(\{\phi_{1},...,\phi_{n}\}\vdash\psi)\Leftrightarrow(\{\phi_{1},...,\phi_{n}\}\Rightarrow\psi).$
Proof.
This result is dependent of (cf. Subsection 3.3): (a) the definition of proof (conditional proof and unconditional proof, cf. Subsection 3.1), (b) the definition of derivable rule (cf. Subsection 3.1), (c) the rule of conditional proof (it is primitive), (d) modus ponens (it is primitive), and (e) rules of commutativity and associativity of conjunction (they are derived above). ∎
With this result, we need not distinguish between a rule $\{\phi_{1},...,\phi_{n}\}\Rightarrow\psi$ and a theorem $\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi$ in proving and using them. Specifically, from a conditional proof $\{\phi_{1},...,\phi_{n}\}\vdash\psi$ or an unconditional proof $\vdash\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi$, we obtain both the rule $\{\phi_{1},...,\phi_{n}\}\Rightarrow\psi$ and the theorem $\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi$; we can use the rule $\{\phi_{1},...,\phi_{n}\}\Rightarrow\psi$ as the theorem $\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi$, and vice versa.
Theorem 7.0.1 (Reflexivity).
$\phi\to\phi.$
Proof.
1 $\phi$ Premise
[…]
Rule 7.0.3 (Transitivity).
$\{\phi\to\psi,\psi\to\chi\}\Rightarrow\phi\to\chi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\psi\to\chi$ Premise
3 $(\psi\to\chi)\to(\phi\to\chi)$ 1: Suffixing
4 $\phi\to\chi$ 2,3: MP
[…]
$\{\phi\vee\psi,\phi\to\chi,\psi\to\chi\}\Rightarrow\chi.$
[…]
Proof.
1 $\phi\vee\psi$ Premise
2 $\phi\to\chi$ Premise
3 $\psi\to\chi$ Premise
4 $\phi\vee\psi\to\chi$ 2,3: “Two-to-one”
5 $\chi$ 1,4: MP
Rule 7.0.5 (Distribution of $\wedge$ over $\vee$, converse).
$(\phi\wedge\psi)\vee(\phi\wedge\chi)\Rightarrow\phi\wedge(\psi\vee\chi).$
Proof.
1 $(\phi\wedge\psi)\vee(\phi\wedge\chi)$ Premise
2     $\phi\wedge\psi$ Assumption
[…]
7 $\phi\wedge\psi\to\phi\wedge(\psi\vee\chi)$ 2–6: CP
8 $\phi\wedge\chi\to\phi\wedge(\psi\vee\chi)$ 2–7: Similarly
9 $\phi\wedge(\psi\vee\chi)$ 1,7,8: Proof by exhaustion
Rule 7.0.6 (Distribution of $\vee$ over $\wedge$).
$\phi\vee(\psi\wedge\chi)\Rightarrow(\phi\vee\psi)\wedge(\phi\vee\chi).$
Proof.
1 $\phi\vee(\psi\wedge\chi)$ Premise
2     $\phi$ Assumption
[…]
6 $\phi\to(\phi\vee\psi)\wedge(\phi\vee\chi)$ 2–5: CP
7     $\psi\wedge\chi$ Assumption
[…]
13 $\psi\wedge\chi\to(\phi\vee\psi)\wedge(\phi\vee\chi)$ 7–12: CP
14 $(\phi\vee\psi)\wedge(\phi\vee\chi)$ 1,6,13: Proof by exhaustion
Rule 7.0.7 (Distribution of $\vee$ over $\wedge$, converse).
$(\phi\vee\psi)\wedge(\phi\vee\chi)\Rightarrow\phi\vee(\psi\wedge\chi).$
Proof.
1 $(\phi\vee\psi)\wedge(\phi\vee\chi)$ Premise
2 $((\phi\vee\psi)\wedge\phi)\vee((\phi\vee\psi)\wedge\chi)$ 1: Distribution (of $\wedge$ over $\vee$, primitive)
3     $(\phi\vee\psi)\wedge\phi$ Assumption
[…]
6 $(\phi\vee\psi)\wedge\phi\to\phi\vee(\psi\wedge\chi)$ 3–5: CP
7     $(\phi\vee\psi)\wedge\chi$ Assumption
8     $(\phi\wedge\chi)\vee(\psi\wedge\chi)$ 7: Distribution (of $\wedge$ over $\vee$, primitive)
9       $\phi\wedge\chi$ Assumption
[…]
12     $\phi\wedge\chi\to\phi\vee(\psi\wedge\chi)$ 9–11: CP
13     $\psi\wedge\chi\to\phi\vee(\psi\wedge\chi)$veeI
14     $\phi\vee(\psi\wedge\chi)$ 8,12,13: Proof by exhaustion
15 $(\phi\vee\psi)\wedge\chi\to\phi\vee(\psi\wedge\chi)$ 7–14: CP
16 $\phi\vee(\psi\wedge\chi)$ 2,6,15: Proof by exhaustion
Rule 7.0.8 (Compatibility of $\to$ with $\wedge$).
$\phi\to\psi\Rightarrow\phi\wedge\chi\to\psi\wedge\chi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\phi\wedge\chi\to\phi$wedgeE
3 $\phi\wedge\chi\to\psi$ 2,1: Transitivity
4 $\phi\wedge\chi\to\chi$wedgeE
5 $\phi\wedge\chi\to\psi\wedge\chi$ 3,4: “One-to-two”
Rule 7.0.9 (Compatibility of $\to$ with $\vee$).
$\phi\to\psi\Rightarrow\phi\vee\chi\to\psi\vee\chi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\psi\to\psi\vee\chi$veeI
3 $\phi\to\psi\vee\chi$ 1,2: Transitivity
4 $\chi\to\psi\vee\chi$veeI
5 $\phi\vee\chi\to\psi\vee\chi$ 3,4: “Two-to-one”
Theorem 7.0.2 ($\top$ is a theorem).
$\top$.⊤ .
Proof.
[…]
2 $\phi\to\phi$ Reflexivity
3 $\top$ 1,2: MP
Theorem 7.0.3 (LEM, theorem form).
$\phi\vee\neg\phi.$
Proof.
[…]
2 $\top$ $\top$ is a theorem
3 $\phi\vee\neg\phi$ 1,2: MP
Theorem 7.0.4 (ECQ, a form without $\bot$).
$\phi\wedge\neg\phi\to\psi.$
Proof.
1 $\phi\wedge\neg\phi$ Premise
[…]
Theorem 7.0.5 (Dual of ECQ, a form without $\top$).
$\phi\to\psi\vee\neg\psi.$
Proof.
1 $\phi$ Premise
[…]
Rule 7.0.10 (DNI).
$\phi\Rightarrow\neg\neg\phi.$
Proof.
1 $\phi$ Premise
2 $\neg\phi\vee\neg\neg\phi$ LEM
[…]
4 $(\phi\wedge\neg\phi)\vee(\phi\wedge\neg\neg\phi)$ 3: Distribution
5 $\phi\wedge\neg\phi\to\neg\neg\phi$ ECQ
6 $\phi\wedge\neg\neg\phi\to\neg\neg\phi$wedgeE
7 $\neg\neg\phi$ 4,5,6: Proof by exhaustion
Rule 7.0.11 (DNE).
$\neg\neg\phi\Rightarrow\phi.$
Proof.
1 $\neg\neg\phi$ Premise
2 $\phi\vee\neg\phi$ LEM
[…]
4 ($(\neg\neg\phi\wedge\phi)\vee(\neg\neg\phi\wedge\neg\phi)$ 3: Distribution
5 $\neg\neg\phi\wedge\phi\to\phi$wedgeE
6 $\neg\neg\phi\wedge\neg\phi\to\phi$ ECQ
7 $\phi$ 4,5,6: Proof by exhaustion
Rule 7.0.12 (Contraposition, negation introduced).
$\phi\to\psi\Rightarrow\neg\psi\to\neg\phi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\neg\psi\to\neg\psi$ Reflexivity
3 $\neg\psi\to\phi\vee\neg\phi$ Dual of ECQ
4 $\neg\psi\to\neg\psi\wedge(\phi\vee\neg\phi)$ 2,3: “One-to-two”
5 $\phi\vee\neg\phi\to\psi\vee\neg\phi$ 1: Compatibility of $\to$ with $\vee$
6 $\neg\psi\wedge(\phi\vee\neg\phi)\to\neg\psi\wedge(\psi\vee\neg\phi)$ 5: Compatibility of $\to$ with $\wedge$
7 $\neg\psi\wedge(\psi\vee\neg\phi)\to(\neg\psi\wedge\psi)\vee(\neg\psi\wedge\neg\phi)$ Distribution
8 $\neg\psi\wedge\psi\to\neg\psi\wedge\neg\phi$ ECQ
9 $\neg\psi\wedge\neg\phi\to\neg\psi\wedge\neg\phi$ Reflexivity
10 ($(\neg\psi\wedge\psi)\vee(\neg\psi\wedge\neg\phi)\to\neg\psi\wedge\neg\phi$ 8,9: “Two-to-one”
11 $\neg\psi\wedge\neg\phi\to\neg\phi$wedgeE
12 $\neg\psi\to\neg\phi$ 4,6,7,10,11: Transitivity
Rule 7.0.13 (Contraposition, negation eliminated).
$\neg\phi\to\neg\psi\Rightarrow\psi\to\phi.$
Proof.
1 $\neg\phi\to\neg\psi$ Premise
2 $\psi\to\neg\neg\psi$ DNI
3 $\neg\neg\psi\to\neg\neg\phi$ 1: Contraposition, negation introduced
4 $\neg\neg\phi\to\phi$ DNE
5 $\psi\to\phi$ 2,3,4: Transitivity
Rule 7.0.14 (Prefixing).
$\phi\to\psi\Rightarrow(\chi\to\phi)\to(\chi\to\psi).$
Proof.
1 $\phi\to\psi$ Premise
2 $(\chi\to\phi)\to(\neg\phi\to\neg\chi)$ Contraposition
3 $\neg\psi\to\neg\phi$ 1: Contraposition
4 ($(\neg\phi\to\neg\chi)\to(\neg\psi\to\neg\chi)$ 3: Suffixing
5 ($(\neg\psi\to\neg\chi)\to(\chi\to\psi)$ Contraposition
6 $(\chi\to\phi)\to(\chi\to\psi)$ 2,4,5: Transitivity
Rule 7.0.15 (RAA, another $\neg$¬I).
$\{\phi\to\psi,\phi\to\neg\psi\}\Rightarrow\neg\phi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\phi\to\neg\psi$ Premise
3 $\neg\psi\to\neg\phi$ 1: Contraposition
4 $\neg\neg\psi\to\neg\phi$ 2: Contraposition
5 $\neg\psi\vee\neg\neg\psi\to\neg\phi$ 3,4: “Two-to-one”
6 $\neg\psi\vee\neg\neg\psi$ LEM
7 $\neg\phi$ 5,6: MP
Rule 7.0.16 (Definition of $\to$ by $\leftrightarrow$).
$\phi\to\psi\Leftrightarrow\phi\wedge\psi\leftrightarrow\phi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\phi\to\phi$ Reflexivity
3 $\phi\to\phi\wedge\psi$ 1,2: “One-to-two”
4 $\phi\wedge\psi\to\phi$wedgeE
[…]
1 $\phi\wedge\psi\leftrightarrow\phi$ Premise
[…]
3 $\phi\wedge\psi\to\psi$wedgeE
4 $\phi\to\psi$ 2,3: Transitivity
Based on some of the above results, it is not difficult to prove the primitive rule $\{\phi\leftrightarrow\psi,\chi\}\Rightarrow\chi[\phi\mapsto\psi]$ of replacement of the Hilbert-style IRL (cf. Subsection 3.2) by induction.
8 Extension to first-order logic
Extending IRL to first-order logic (FOL) is similar to what is done classically.
Let $\{x,y,...\}$, $\{s,t,...\}$, $\{P,Q,...\}$ and $\{\phi,\psi,...\}$ be the FOL variables, terms, predicates and formulas, respectively.
Suppose that in a substitution $\phi\{t/x\}$ or a replacement $\phi[s\mapsto t]$, $s$ and $t$ are not bound variables in $\phi$ (if they are, rename them before substitution or replacement).
The extension of the propositional logic IRL to first-order logic is defined as a FOL that contains:
  1. 1.
    All theorems and rules of IRL where propositional formulas are extended to be any FOL formulas.
  2. 2.
    Axioms for equality:
    • Reflexivity: $s=s$. Symmetry and transitivity can be derived.
    • Replacement: $(s=t)\wedge\phi\to\phi[s\mapsto t]$.
  3. 3.
    Axioms for quantifiers:
    • Universal generalization: $\phi(y)\to\forall x(\phi(x))$ if $y$ is arbitrary.
    • Universal instantiation: $\forall x(\phi(x))\to\phi\{t/x\}$ where $t$ can be any term.
    • “One-to-many”: $\forall x(\phi\to\psi(x))\to(\phi\to\forall x(\psi(x)))$ where $x$ does not occur free in $\phi$. Its converse can be derived.
    • Duality: $\neg\forall x(\phi)\to\exists x(\neg\phi)$. Another one $\neg\exists x(\phi)\to\forall x(\neg\phi)$ can be derived.
    • Existential instantiation: $\exists x(\phi(x))\to\phi\{t_{0}/x\}$ where $t_{0}$ is unknown, so it must not occur earlier in the proof and it also must not occur in the conclusion of the proof.
    • Existential generalization (redundant): $\phi\{t_{0}/x\}\to\exists x(\phi(x))$ where $t_{0}$ is some (known) term.
    • “Many-to-one” (redundant): $\forall x(\phi(x)\to\psi)\to(\exists x(\phi(x))\to\psi))$ where $x$ does not occur free in $\psi$. Its converse can be derived so is redundant as well.
The following are some derived results.
Theorem 8.0.1 (“One-to-many”, converse).
$(\phi\to\forall x(\psi(x)))\to\forall x(\phi\to\psi(x))$
where $x$ does not occur free in $\phi$.
Proof.
1 $\phi\to\forall x(\psi(x))$ Premise
2 $\forall x(\psi(x))\to\psi(y)$ Universal instantiation ($y$ is arbitrary)
3 $\phi\to\psi(y)$ 1,2: Transitivity
4 $\forall x(\phi\to\psi(x))$ 3: Universal generalization (as $y$ is arbitrary from line 2)
Theorem 8.0.2 (Duality, existential to universal quantifier).
$\neg\exists x(\phi)\to\forall x(\neg\phi).$
Proof.
1 $\neg\exists x(\phi)$ Premise
2 $\neg\forall x(\neg\phi)\to\exists x(\neg\neg\phi)$ Duality, universal to existential quantifier
3 $\neg\forall x(\neg\phi)\to\exists x(\phi)$ 2: Double negation
4 $\neg\neg\forall x(\neg\phi)$ 1,3: Modus tollens
5 $\forall x(\neg\phi)$ 4: Double negation
Theorem 8.0.3 (Existential generalization).
$\phi\{t_{0}/x\}\to\exists x(\phi(x))$
where $t_{0}$ is some (known) term.
Proof.
1 $\phi\{t_{0}/x\}$ Premise
2 $\forall x(\neg\phi(x))\to\neg\phi\{t_{0}/x\}$ Universal instantiation
3 $\neg\forall x(\neg\phi(x))$ 1,2: Modus tollens
4 $\exists x(\neg\neg\phi(x))$ 3: Duality
5 $\exists x(\phi(x))$ 4: Double negation
Theorem 8.0.4 (“Many-to-one”).
$\forall x(\phi(x)\to\psi)\to(\exists x(\phi(x))\to\psi))$
where $x$ does not occur free in $\psi$.
Proof.
1 $\forall x(\phi(x)\to\psi)$ Premise
2 $\phi(y)\to\psi$ 1: Universal instantiation ($y$ is arbitrary)
3 $\neg\psi\to\neg\phi(y)$ 2: Contraposition
4 $\forall x(\neg\psi\to\neg\phi(x))$ 3: Universal generalization (as $y$ is arbitrary from line 2)
5 $\neg\psi\to\forall x(\neg\phi(x))$ 4: “One-to-many”
6 $\neg\forall x(\neg\phi(x))\to\psi$ 5: Contraposition, Double negation
7 $\exists x(\phi(x))\to\psi$ 6: Duality
Theorem 8.0.5 (“Many-to-one”, converse).
$(\exists x(\phi(x))\to\psi)\to\forall x(\phi(x)\to\psi)$
where $x$ does not occur free in $\psi$.
Proof.
1 $\exists x(\phi(x))\to\psi$ Premise
2 $\neg\psi\to\neg\exists x(\phi(x))$ 1: Contraposition
3 $\neg\psi\to\forall x(\neg\phi(x))$ 2: Duality
4 $\forall x(\neg\psi\to\neg\phi(x))$ 3: “One-to-many”, converse
5 $\neg\psi\to\neg\phi(y)$ 4: Universal instantiation ($y$ is arbitrary)
6 $\phi(y)\to\psi$ 5: Contraposition
7 $\forall x(\phi(x)\to\psi)$ 6: Universal generalization (as $y$ is arbitrary)
Theorem 8.0.6 (Universal conjunction).
$\forall x(\phi(x)\wedge\psi(x))\leftrightarrow\forall x(\phi(x)\wedge\forall x(\psi(x).$
Proof.
1 $\forall x(\phi(x)\wedge\psi(x))$ Premise
2 $\phi(y)\wedge\psi(y)$ 1: Universal instantiation ($y$ is arbitrary)
3 $\phi(y)$ 2: Conjunction elimination
4 $\forall x(\phi(x))$ 3: Universal generalization (as $y$ is arbitrary)
5 $\forall x(\psi(x))$ 2–4: Similarly
6 $\forall x(\phi(x)\wedge\forall x(\psi(x)$ 4,5: Conjunction introduction
The converse can be proved in the reverse order. ∎
Theorem 8.0.7 (Existential disjunction).
$\exists x(\phi(x)\vee\psi(x))\leftrightarrow\exists x(\phi(x)\vee\exists x(\psi(x)).$
Proof.
1 $\exists x(\phi(x)\vee\psi(x))$ Premise
2 $\phi(a)\vee\psi(a)$ 1: Existential instantiation
3 $\phi(a)\to\exists x(\phi(x))$ 2: Existential generalization
4 $\phi(a)\to\exists x(\phi(x))\vee\exists x(\psi(x))$ 3: Weakening
5 $\psi(a)\to\exists x(\phi(x))\vee\exists x(\psi(x))$ 2–4: Similarly
6 $\exists x(\phi(x))\vee\exists x(\psi(x))$ 2,4,5: Proof by exhaustion
1 $\exists x(\phi(x))\vee\exists x(\psi(x))$ Premise
2 $\exists x(\phi(x))\to\phi(a)$ Existential instantiation
3 $\exists x(\phi(x))\to\phi(a)\vee\psi(a)$ 2: Weakening
4 $\phi(a)\vee\psi(a)\to\exists x(\phi(x)\vee\psi(x))$ Existential generalization
5 $\exists x(\phi(x))\to\exists x(\phi(x)\vee\psi(x))$ 3–4: Transitivity
6 $\exists x(\psi(x))\to\exists x(\phi(x)\vee\psi(x))$ 2–5: Similarly
7 $\exists x(\phi(x)\vee\psi(x))$ 1,5,6: Proof by exhaustion
Theorem 8.0.8 (Quantifier switching).
$\exists x\forall y(\phi(x,y))\to\forall y\exists x(\phi(x,y)).$
Proof.
1     $\exists x\forall y(\phi(x,y))$ Assumption
2     $\neg\forall y\exists x(\phi(x,y))$ Assumption
3     $\forall y(\phi(a,y))$ 1: Existential instantiation
4     $\exists y\forall x(\neg\phi(x,y))$ 2: Duality
5     $\forall x(\neg\phi(x,b))$ 4: Existential instantiation
6     $\phi(a,b)$ 3: Universal instantiation
7     $\neg\phi(a,b)$ 5: Universal instantiation
8     $\bot$ 6,7: LNC
9 $\exists x\forall y(\phi(x,y))\wedge\neg\forall y\exists x(\phi(x,y))\to\bot$ 1–8: Conditional proof
10 $\exists x\forall y(\phi(x,y))\to\forall y\exists x(\phi(x,y))$ 9: Implication and inconsistency
Meta Rule 8.0.1 (Equivalence of implication theorem and disjunction theorem in FOL).
$(\vdash\forall x\,(\phi(x)\to\psi(x)))\Leftrightarrow(\vdash\forall x\,(\neg\phi(x)\vee\psi(x))).$
Proof.
1 $\vdash\forall x\,(\phi(x)\to\psi(x))$ Premise
2     $\forall x\,(\phi(x)\to\psi(x))$ 1 (assumed theorem)
3     $\phi(y)\to\psi(y)$ 2: Universal instantiation ($y$ is arbitrary)
4 $\vdash\phi(y)\to\psi(y)$ 2–3: Unconditional proof
5 $\vdash\neg\phi(y)\vee\psi(y)$ 4: Meta Rule 3.4.1 of IRL
6     $\neg\phi(y)\vee\psi(y)$ 5 (theorem indicated in an outer proof)
7     $\forall x\,(\neg\phi(x)\vee\psi(x))$ 6: Universal generalization (as $y$ is arbitrary)
8 $\vdash\forall x\,(\neg\phi(x)\vee\psi(x))$ 6–7: Unconditional proof
The converse can be proved in the reverse order. ∎
Bounded quantifiers
It is convenient if including bounded quantifiers or restricted quantifiers in a first-order language.
Definition 8.0.1 (Bounded quantifiers).
Let $\chi$ be a unary formula, $\phi$ be any formula. Let $X=\{x\mid\chi(x)\}$. Then $x\in X$ if and only if $\chi(x)$. The bounded quantifiers $\exists x\in X$ and $\forall x\in X$ are defined by the rules:
  • $\exists x\in X\,(\phi)\Leftrightarrow\exists x\,(\chi(x)\wedge\phi)$.
  • $\forall x\in X\,(\phi)=\neg\exists x\in X\,(\neg\phi)\Leftrightarrow\neg\exists x\,(\chi(x)\wedge\neg\phi)\Leftrightarrow\forall x\,\neg(\chi(x)\wedge\neg\phi)$, i.e., $\forall x\in X\,(\phi)\Leftrightarrow\forall x\,(\neg\chi(x)\vee\phi)$.
Remark 8.0.1.
The universal bounded quantifier $\forall x\in X$ can not be defined by $\forall x\in X\,(\phi)\Leftrightarrow\forall x\,(\chi(x)\to\phi)$ since $\chi(x)\to\phi$ and $\neg\chi(x)\vee\phi$ are not equivalent in IRL – however, if $\forall x\in X\,(\phi)$ represents a theorem, the two expressions are equivalent according to the above Meta Rule 8.0.1.
9 Discussion
Normal situation
An arbitrary proposition can be tautological, contradictory, or contingent. Most generally, there are possibly infinitely many T-set values for a contingency ($10<x<$ 1) whereas there is only one value for a tautology ($x=1$) and one value for a contradiction ($x=0$). This is just similar to the unit interval $[0,1]$ (no matter it is different as totally ordered) where the number of interior points is uncountable while the number of boundary points is only two. In this sense, it is the normal situation for a proposition being contingent rather than tautological or contradictory. Subsequently in this section, when saying “in the normal situation” we mean propositions mentioned are all contingent.
Interpretation of Generalized ECQ of implication
In IRL, the Generalized ECQ $\psi\to(\phi\to\psi)$ does not hold while the Generalized ECQ of implication $(\phi\to\psi)\to(\chi\to(\phi\to\psi))$ does.
Generalized ECQ $\psi\to(\phi\to\psi)$ means that, for any propositions $\phi$ and $\psi$, there is always an implication from $\phi$ to $\psi$ in the normal situation (where $\psi$ can be true so $\phi\to\psi$ can be true). This is not consistent with the original meaning of implication as it is not ensured for a “certain mechanism” of implication in this situation.
Whereas Generalized ECQ of implication $(\phi\to\psi)\to(\chi\to(\phi\to\psi))$ means that if there exists the “certain mechanism” of implication from $\phi$ to $\psi$, then this mechanism is irrelevant to any proposition $\chi$. This does not conflict with the original meaning of implication at all.
The “worst” case for Generalized ECQ of implication is when $\chi=\neg(\phi\to\psi)$. From IRL theorems ($(\neg\phi\to\phi)\leftrightarrow(\top\to\phi)$ and $(\top\to\phi)\to\phi$, we have $(\top\to(\phi\to\psi))\leftrightarrow(\phi\to\psi)$ in the worst case. This can be understood as a way to interpret $\top\to(\phi\to\psi)$ as $\phi\to\psi$. In contrast, since the general form $(\top\to\phi)\leftrightarrow\phi$ does not hold (as $\phi\to(\top\to\phi)$ does not hold), $\top\to\phi$ can not be interpreted as $\phi$ generally.
Interpretation of ECQ
A simple conditional proof of ECQ $\phi\wedge\neg\phi\to\psi$:
1 $\phi$ Premise 2 $\neg\phi$ Premise 3 $\phi\vee\psi$ 1: Disjunction introduction 4 $\psi$ 2,3: Disjunctive syllogism
Hence ECQ is necessary for disjunction introduction and disjunctive syllogism which do not conflict to the original meaning of implication. Therefore, ECQ contributes to the goal of strength for the system (cf. Subsection 2.6).
Furthermore, ECQ $\bot\to\phi$ and its dual ϕitalic-$\phi\to\top$ can be interpreted to intuition as “if an always-false thing is (was) true, then anything is (would be) true” and “in any cases, an always-true thing is true”, respectively, as similarly mentioned, e.g., in (Ceniza, 1988).
ECQ is an abnormal situation, for which a “standard (set of rules for) interpretation” may be explored for practical purpose. For example, given a conditional $P\to Q$: if $P$ is a contradiction (e.g. it is well known that $P$ is false), then $P\to Q$ can be interpreted as “if the false thing like $P$ were true, anything including $Q$ should be true”; if $Q$ is a tautology (e.g. it is a well known fact), then $P\to Q$ can be interpreted as “in any cases including $P$, an always true thing like $Q$ is true”.
10 Concluding remarks
By the original meaning of implication, the formula $\neg\phi\vee\psi\to(\phi\to\psi)$ is generally invalid, so the material implication $(\phi\to\psi)\leftrightarrow\neg\phi\vee\psi$ is not valid in general. Defining implication according to its original meaning is the key to avoid invalid results. The logic IRL of this work has the expected property that the invalid class of $\neg\phi\vee\psi\to(\phi\to\psi)$ is removed while fundamental laws like LEM and double negation, LNC and ECQ, conjunction elimination and disjunction introduction, and transitivity of implication and disjunctive syllogism, are all retained. Based on information provided with IRL, more points are summarized as follows.
  1. 1.
[…]
  2. 2.
    Truth value of implication. Since implication is not truth functional, valuation of $\phi\to\psi$ can not be generally done by valuation of $\phi$ and $\psi$, but it can be “estimated” via the procedure: (a) $\phi\to\psi$ is a tautology, if and only if $\neg\phi\vee\psi$ is a tautology; (b) $\phi\to\psi$ is a contradiction, if $\neg\phi\vee\psi$ is a contradiction (i.e. $\phi$ is a tautology and $\psi$ is a contradiction); (c) the T-set of $\phi\to\psi$ is a subset (possibly empty) of the T-set of $\neg\phi\vee\psi$, if $\neg\phi\vee\psi$ is a contingency. T-sets of primitive implications, like any primitive propositions, are given (e.g. by physical mechanisms) rather than derived.
  3. 3.
    Classical logic is sound and complete with regard to its own semantics. In the sense that it contains theorems that are not generally valid with respect to semantics based on the original meaning of implication, classical logic is less “sound” than “complete” (actually it is post-complete), while some weaker systems are in the opposite.
  4. 4.
    The logic IRL of this work, in which the generally invalid classical theorems, including especially the class of $\neg\phi\vee\psi\to(\phi\to\psi)$, are removed, should be used instead of classical logic for the general case (the system contains any types of propositions including especially contingencies). IRL and classical logic are the same if and only if there is no (primitive) contingent propositions in a system.
  5. 5.
    An interpreted formal system has no contingent statements. Precisely, for a system under a given interpretation, every sentence is simply either true or false, no longer relevant to “tautological, contradictory or contingent”. It is common that a theory has an intended interpretation or a standard model or usual model. For example, Tarski’s axioms for Euclidean geometry has its standard model of Euclidean plane geometry, and the standard model for Peano axioms and Robinson arithmetic is the natural numbers. In such contexts when a system contains no contingencies, classical logic can be used safely.